### v3

# Quantum algorithms for calculating determinant and inverse of matrix and solving linear algebraic systems

Abstract
We propose quantum algorithms for calculating the determinant and inverse of an $(N-1)\times(N-1)$ matrix (depth is $O(N^{2}\log N)$) which is a simple modification of the algorithm for calculating the determinant of an $N\times N$ matrix (depth is $O(N\log^{2}N)$. The basic idea is to encode each row of the matrix into a pure state of some quantum system. In addition, we use the representation of the elements of the inverse matrix in terms of algebraic complements. Combination of this algorithm with the quantum algorithm for matrix multiplication, proposed in our previous work, yields the algorithm for solving systems of linear algebraic equations (depth is $O(N\log^{2}N)$. Measurement of the ancilla state with output 1 (probability is $\sim 2^{-O(N\log N)}$) removes the garbage acquired during calculation. Appropriate circuits for all three algorithm are presented and have the same estimation $O(N\log N)$ for the size.
Keywords: quantum algorithm, quantum circuit, determinant, inverse matrix, linear system, depth of algorithm, quantum measurement
I Introduction
Quantum algorithms that exploit principles of quantum physics can be run on quantum computers and promise to offer considerable advantages over classical counterparts. Algorithms for quantum calculations are an increasingly prominent field in quantum information research. Notable examples include Shor’s factoring algorithm Sho1 ; Sho2 , Deutch parallelism Deu , Grover’s search algorithm Gro , Quantum Fourier transform QFT1 ; QFT2 ; QFT3 and Quantum phase estimation Wang ; QFT3 . Further advancements include the HHL algorithm for solving systems of linear equations HHL ; HHL1 ; HHL2 ; HHL3 ; HHL4 ; HHL5 ; HHL6 ; HHL7 , in which the crucial point is the conditional rotation of the state of the one-qubit ancilla. This operation, where inversion of the eigenvalue requires involving classical computing, is placed between the blocks of the direct and inverse phase estimation algorithms. HHL algorithm exponentiates a Hermitian matrix of the considered linear system, which can be performed via the Trotterization formula T ; AT . There is extensive literature devoted to improvement of this algorithm, especially its scale and complexity BHMT ; CKS ; GSLW ; LC ; LYC ; AL ; JRS ; SSO ; AlL ; TAWL . However, the principal problem of eigenvalue inversion is not resolved via quantum methods yet. For this purpose, for instance, the quantum state tomography can be involved to obtain the eigenvalues of the exponent of Hermitian operator (the matrix of the linear system) T . Nevertheless, the HHL algorithm combining the phase estimation and classical inversion of eigenvalues is very popular in realization of various algorithms of quantum computation and it is widely applicable in quantum machine learning A ; HHL2 ; RML . In particular, it is applied in the least-square linear-regression algorithms WBL ; SSP ; Wang2 working with large data sets, in quantum algorithms for singular-value decomposition GSLW ; G , in algorithms for solving linear B and nonlinear LKKLTC differential equations.
We also have to refer to algorithms for matrix manipulations based on the Trotterization method and the Baker-Champbell-Housdorff formula for exponentiating matrices ZhaoL1 . In GSLW , algebraic manipulations with matrices are implemented using unitary transformations that include the matrices as specific blocks (block-encoding model). A similar block-encoding approach was used in the algorithm of DFZ_2020 to embed the inverse of the matrix of an algebraic system into the unitary transformation. In LWWZ , an algorithm for matrix multiplication using binary encoding of matrix elements in the states of quantum systems followed by binary multiplication using Toffoli gates was proposed. Matrix multiplication over rings, including the multiplication of Boolean matrices was studied in KN .
In our recent work QZKW_arxive2022 , we presented alternative algorithms for matrix manipulations, such as sum and product of two matrices, calculation of determinant, matrix inversion, and solving linear systems, based on special unitary transformations of a quantum system. The main distinction of these algorithms from those mentioned above is the encoding of matrix elements into the pure states of specific (sub)systems as the probability amplitudes. However, the realization of those unitary operators in terms of the basic quantum operations was not explored in that study.
Later, in ZQKW_arxiv2023 , we proposed the implementation of unitary transformations corresponding to vector inner product, matrix sum, and product using multi-qubit conditional operators (reducible to Toffoli gates) and Hadamard transformations, including the ancilla measurement for garbage removal. The depth or complexity of these algorithms increases logarithmically with the matrix dimension, or even remains independent of it (in the case of matrix sum), demonstrating the advantage of these algorithms compared to their classical counterparts.
In this paper, we propose the algorithm for matrix inversion with its application to solving systems of linear equations. The principal part of this algorithm is the algorithm for computing the determinant of a square matrix. It is built on the ideas of Refs. QZKW_arxive2022 ; ZQKW_arxiv2023 :
    Encode matrix elements into the states of a certain quantum subsystems through probability amplitudes in such a way that all necessary multiplications of matrix elements are performed automatically and appear in the probability amplitudes of the initial state of the whole system;
    Use multiqubit control-SWAP and control-Z operators together with ancilla to organize proper signs of each term in the superposition state and collect them in certain state of the considered system;
    Use ancilla measurement to remove the garbage and present the desirable result as a pure state of a certain quantum system. Normalization of this state is defined at the stage of ancilla measurement.
We follow the above steps to propose an algorithm for calculating the determinant of a square matrix whose $N$ rows are encoded into the pure states of $N$ separate quantum subsystems. Notice that creating such initial state is a special problem by itself, but we do not handle this problem here. The state of the whole system results in all possible products of $N$ elements from different rows. One only needs to select the proper products to calculate the determinant. We propose such an algorithm and construct the quantum circuit realizing it. While developing the algorithm, our focus is on the sizes and depths of its specific blocks. Essentially, our algorithm is a quantum realization of the cofactor expansion classical algorithm, whose depth exponentially increases with matrix dimension. However, the resulting depth of our algorithm is $O(N\log^{2}N)$ (or $O(N^{2}\log N)$ in the worst case) with the size $O(N\log N)$. The final step of this algorithm is the one-qubit ancilla measurement, which removes all the garbage. The probability of obtaining the required result of the ancilla measurement (which is 1) exponentially decreases with the matrix dimension $N$, which requires exponentially increasing number of evaluations of the algorithm. This is a central problem that we plan to overcome in further work.
It is remarkable that the minor modification of the algorithm for calculating the determinant yields an alternative quantum algorithm for matrix inversion using its definition in terms of algebraic complements. We again use the row-wise encoding of the matrix elements into the state of a quantum system so that all necessary products of matrix elements appear automatically in the superposition state, thus one only needs to properly sort these products. The depth of this algorithm is $O(N^{2}\log N)$ (or $O(N^{2}\log^{2}N)$ in the worst case) and the size is $O(N\log N)$. Based on such inversion algorithm we develop an alternative algorithm for solving systems of linear algebraic equations. Thus, we do not use Trotterization methods and classical subroutine for inversion of eigenvalues. Therefore the proposed algorithm is completely quantum and it is an alternative to the inversion algorithm considered in HHL ; T ; TAWL . The estimations for the depth and size of this algorithm remain the same.
In relation with the question of algorithm complexity, we shall mention that the special methods of classical manipulations with matrices allow to reduce the depth of the algorithms for determinant calculation to $O(\log^{2}N)$ Ber ; BAEM with the size $O(N^{3.496})$. In our case we use the Leibniz formula as a definition of the matrix determinant and definition of the inverse matrix in terms of the algebraic complements. Our quantum algorithms based on these definitions obviously reveal advantages in comparison with classical realizations of these definitions. It is likely that using specific classical algorithms for matrix manipulations and developing their quantum analogues lead to further improvement in the depth and size of quantum algorithms calculating the determinant and inverse matrix. This issue is not considered in our paper. The main novelty is that we are promoting the method for representation of the matrix elements as probability amplitudes of pure states and reveal profits out of such representation in application to the standard algorithms for matrix manipulations.
The paper is organized as follows. The algorithm for calculating the matrix determinant is proposed in Sec.II. This algorithm is taken as a basis for the algorithm calculating the inverse matrix presented in Sec. III. Based on that results, the algorithm for solving systems of linear algebraic equations is proposed in Sec. IV. The main results are discussed in Sec.V. Appendix, Sec.VI, contains the modified algorithms for matrix multiplication.
II Determinant
To calculate the determinant of $N\times N$ matrix $\tilde{A}$ (we assume $N=2^{n}$) with elements $a_{jk}$ ($j,k=0,1,\cdots,N-1$),
$\tilde{A}=\left(\begin{array}[]{cccc}a_{00}&a_{01}&\cdots&a_{0(N-1)}\cr a_{10}&a_{11}&\cdots&a_{1(N-1)}\cr\cdots&\cdots&\cdots&\cdots\cr a_{(N-2)0}&a_{(N-2)1}&\cdots&a_{(N-2)(N-1)}\cr a_{(N-1)0}&a_{(N-1)1}&\cdots&a_{(N-1)(N-1)}\end{array}\right),$ (6)
we encode each row of $N$ elements into the pure state $|\Psi_{j}\rangle$ of the $n$-qubit subsystem $S_{j}$,
$|\Psi_{j}\rangle=\sum_{k=0}^{N-1}a_{jk}|k\rangle_{S_{j}},\;\;j=0,1,\dots,N-1.$ (7)
Representing rows of the matrix $\tilde{A}$ in the form of pure states (7) presumes $N$ obvious constraints for the matrix elements $a_{ij}$ predicted by the normalization of pure states:
$\sum_{k=0}^{N-1}|a_{jk}|^{2}=1,\;\;j=0,\dots,N-1.$ (8)
These constraints can be smoothened by incorporating additional terms with different probability amplitudes into the states $|\Psi_{i}\rangle$, but this issue will be addressed in future study. The whole system is the composition of the subsystems, $S=\bigcup_{j}S_{j}$, and the pure state of the whole system reads
$|\Phi_{0}\rangle$ $=$ $|\Psi_{0}\rangle\cdots|\Psi_{N-1}\rangle$ (9)
$=$ $\sum_{k_{0},\dots,k_{N-1}=0}^{N-1}a_{0k_{0}}\dots a_{(N-1)k_{N-1}}|k_{0}\rangle_{S_{0}}\cdots|k_{N-1}\rangle_{S_{N-1}},$
which includes $N^{(S)}=nN$ qubits. In computing the determinant, we need only those terms in which all $k_{j}$ are different. We select these terms from the whole pure state of our system:
$|\Phi_{0}^{(D)}\rangle$ $=$ $\sum_{k_{0}\neq\dots\neq k_{N-1}}a_{0k_{0}}\dots a_{(N-1)k_{N-1}}|k_{0}\rangle_{S_{0}}\cdots|k_{N-1}\rangle_{S_{N-1}}+|g_{1}\rangle,$ (10)
where $|g_{1}\rangle$ collects all extra terms which form garbage to be removed later.
We have to arrange the appropriate signs for each term of the first part in the right-hand side of (10). This can be achieved by setting the subsystem $S_{j}$ (in each term) into the state $|j\rangle$ using the SWAP operators together with sign-switching. This pair of operators performs permutation properly altering the sign of the term.
Thus, our purpose is to exchange the states of subsystems $S_{j}$ ($j=0,\dots,N-1$) so that the state of the subsystem $S_{k}$ becomes $|k\rangle_{S_{k}}$ in each term of the first part of superposition state (10). For that we introduce the projectors $P_{S_{j}}^{(k)}=|k\rangle_{S_{j}}\;{}_{S_{j}}\langle k|$, ($k=0,\dots N-2,$ $j=k+1,\dots N-1$) and the $\tilde{n}_{k}$-qubit ancillae $A_{k}$, ($\tilde{n}_{k}=\lceil\log(N-k)\rceil$, $k=0,\dots N-2$). The ancilla $A_{k}$ serves to save the index $j$ of the subsystem $S_{j}$ whose state is $|k\rangle$ with $j>k$. Therefore, the number of qubits $\tilde{n}_{k}$ in $A_{k}$ is less then $n=\log N$ unless $k=0$: $\tilde{n}_{k}\leq n$. For instance, in Fig. 2 for the case $N=4$, there are three ancillae $A_{0},A_{1},A_{2}$. Ancilla $A_{2}$ has to encode either only one state $|3\rangle$ or nothing, therefore one qubit is enough: we can encode $|3\rangle$ by the state $|1\rangle$ of ancilla $A_{2}$. Next, $A_{1}$ has to encode two states $|2\rangle$ and $|3\rangle$ or nothing. Two qubits are enough for that: $|2\rangle\rightarrow|1\rangle,|3\rangle\rightarrow|2\rangle$. Similarly, $A_{0}$ requires 2 qubits as well. Using the projectors $P_{S_{j}}^{(k)}$ and ancillae $A_{k}$ we can prepare the operators $V_{S_{j}A_{k}}$, which relates the index $j$ of $S_{j}$ to the appropriate state $|\tilde{j}\rangle_{A_{k}}$ of the ancilla $A_{k}$:
$V_{S_{j}A_{k}}$ $=$ $P_{S_{j}}^{(k)}\otimes Z_{k}(j)\sigma^{(x)}_{kj}+(I_{S_{j}}-P_{S_{j}}^{(k)})\otimes I_{A_{k}},$ (11)
where $\sigma^{(x)}_{kj}$ includes the set of $\sigma^{(x)}$-operators applied to the ground state of the ancilla $A_{k}$ to form the state $|\tilde{j}\rangle_{A_{k}}$; $Z_{k}(j)$ changes the sign of the ancilla state $|\tilde{j}\rangle_{A_{k}}$ (it counts permutations and can be represented by either the $\sigma^{(z)}$-operator applied to one of the excited qubits, or by the $\sigma^{(x)}\sigma^{(z)}\sigma^{(x)}$-operator applied to one of the ground-state qubits (only to one!) of $A_{k}$). To arrange the exchange of states between $S_{j}$ and $S_{k}$, we introduce the projectors $P_{A_{k}}^{(j)}=|\tilde{j}\rangle_{A_{k}}\;{}_{A_{k}}\langle\tilde{j}|,$ $k=0,\dots N-2,$ $j=k+1,\dots N-1,$ $\tilde{j}=0,\dots N-k-2,$ and construct the operators $U_{S_{k}S_{j}}$:
$U_{S_{k}S_{j}}$ $=$ $\Big{(}SWAP_{S_{k},S_{j}}\otimes P_{A_{k}}^{(j)}$ (12)
$+$ $I_{S_{k}S_{j}}\otimes(I_{A_{k}}-P_{A_{k}}^{(j)})\Big{)},\;\;k\neq j.$
The circuits for the operators $V_{S_{j}A_{k}}$ and $U_{S_{k}S_{j}}^{(k)}$ are given in Fig.1(a) (the general matrix case) and in Fig.1(b) (the particular $4\times 4$-matrix case).
Figure 1: $V$-and $U$-operators: (a) the general case; (b) the $4\times 4$-matrix case, then $\tilde{n}_{0}=2$, $\tilde{n}_{1}=\tilde{n}_{2}=1$, $\tilde{n}_{k}<n$, therefore, $A_{1}$ is a one-qubit ancilla.
The depth of each operator $V_{S_{j}A_{k}}$ is $O(n)$ (due to the projectors $P_{S_{j}}^{(k)}$ with $n$ control qubits and we assume that all $\sigma^{(x)}$-operators of $\sigma^{(x)}_{kj}$ are applied in parallel to different qubits of $a_{k}$), while the depth of $U_{S_{k}S_{j}}$ is the of depth of the control-SWAP-operator which includes $n$ swaps each controlled by $n$ qubits, i.e. the total depth is $n\cdot n=n^{2}$ (or $O(n^{2})$).
We prepare the operators
$W_{k}=\prod_{j=k+1}^{N-1}U_{S_{k}S_{j}}V_{S_{j}A_{k}}.$ (13)
Applying $W_{k}$ to the state $|\Phi_{0}\rangle|0\rangle_{A}$ we put $S_{k}$ into the state $|k\rangle_{S_{k}}$ in each term of the first part of (10). Although $W_{k}$ has product of $N-k-1$ pairs of $U$- and $V$-operators, only one pair of such operators is applied to each particular term in $|\Phi_{0}\rangle$ which is due to the structure of the projector operators $P^{(k)}_{S_{j}}$. We assume that this means the $UV$-products can be applied in parallel and thus the depth of $W_{k}$ is $O(n)+O(n^{2})=O(n^{2})$. Otherwise we have to multiply the obtained depth by $O(N)$. The operator $\prod_{k=0}^{N-2}W_{k}$ provides the needed sign of the probability amplitudes which will compose the expression for the determinant. To collect all of them in a single probability amplitude we apply the Hadamard transformation to each qubit of ancillae $A_{k}$, $k=0,\dots,N-2$, and denote this transformations $H_{A}$. There are $\tilde{N}$ Hadamard operators in $H_{A}$,
$\tilde{N}=\sum_{k=0}^{N-2}\tilde{n}_{k}=O(N\log N).$ (14)
Applying the operator $W^{(1)}_{SA}$,
$W^{(1)}_{SA}=H_{A}\prod_{k=0}^{N-2}W_{k},$ (15)
to the state $|\Phi_{0}\rangle|0\rangle_{A}$, $A=\bigcup_{k=0}^{N-2}A_{k}$, and selecting the terms which carry the information about the determinant (these terms correspond to the ground state of $A$) we obtain
$|\Phi_{1}^{(D)}\rangle=W^{(1)}_{SA}|\Phi_{0}\rangle|0\rangle_{A}$ (16)
$=\frac{\det(A)}{2^{\tilde{N}/2}}\;|0\rangle_{S_{0}}|1\rangle_{S_{1}}\otimes|{N-1}\rangle_{S_{N-1}}|0\rangle_{A}+|g_{2}\rangle_{SA},$
where
$\det(A)$ $=$ $\sum_{k_{0}\neq\dots\neq k_{N-1}}\epsilon_{k_{0}\dots k_{N-1}}a_{k_{0}0}\dots a_{k_{N-1}(N-1)},$ (17)
and $\epsilon_{k_{0}\dots k_{N-1}}$ is the permutation symbol. The depth of $W^{(1)}$ is $O(Nn^{2})$. Next, we can set states of all $S_{j}$ in the first term of $|\Phi_{1}\rangle$ into the $|0\rangle_{S_{j}}$ states. To this end we apply the $\sigma^{x}$-operators to the excited qubits of $S$, i.e. prepare the operator $X_{S}$,
$X_{S}=\prod_{j=1}^{N-1}X_{S_{j}}(j),$ (18)
with $X_{S_{j}}(j)$ being the set of $\sigma^{(x)}$ operators acting on the excited qubits of $S_{j}$ in the state $|j\rangle_{S_{j}}$. The depth of $X_{S}$ is $O(1)$:
$|\Phi_{2}^{(D)}\rangle=X_{S}|\Phi_{1}^{(D)}\rangle$ (19)
$=\frac{\det(A)}{2^{\tilde{N}/2}}\;|0\rangle_{S}|0\rangle_{A}+|g_{3}\rangle_{SA}.$
We notice that the operators $H_{A}$ in $W^{(1)}_{SA}$ and operator $X_{S}$ can be applied simultaneously because they effect on different qubits. We separate them to underline the distinction in their functions.
Next, we remove the garbage. For that we introduce the projector $P_{SA}=|0\rangle_{S}|0\rangle_{A}\;{}_{A}\langle 0|_{S}\langle 0|$, include the one-qubit ancilla $B$ in the ground state and construct the operator $W^{(2)}_{SAB}$,
$W^{(2)}_{SAB}=P_{SA}\sigma^{(x)}_{B}+(I_{SA}-P_{SA})I_{B},$ (20)
whose depth is $O(Nn+Nn)=O(Nn)$. Applying it to $|\Phi_{2}^{(D)}\rangle|0\rangle_{B}$ we obtain
$|\Phi_{3}^{(D)}\rangle=W^{(2)}_{SAB}|\Phi_{2}^{(D)}\rangle|0\rangle_{B}$ (21)
$=\frac{\det(A)}{2^{\tilde{N}/2}}\;|0\rangle_{S}|0\rangle_{A}|1\rangle_{B}+|g_{3}\rangle_{SA}|0\rangle_{B}.$
Now we measure the ancilla $B$ with the output $|1\rangle_{B}$. The probability of this result of measurement is $2^{-\tilde{N}}|{\det(A)}|^{2}$. After the measurement we get the following state:
$|\Phi_{4}^{(D)}\rangle=|\Psi^{(D)}_{out}\rangle|0\rangle_{A},\;\;|\Psi^{(D)}_{out}\rangle=\arg(\det(A))\;|0\rangle_{S},$ (22)
The particular circuit for calculating the $4\times 4$ determinant using the proposed algorithm is given in Fig. 2. The circuit for the determinant of $N\times N$ matrix includes $N\log N$ qubits of the subsystem $S$, $\tilde{N}=O(N\log N)$ qubits of the subsystem $A$ and one qubit of the ancilla $B$, therefore the size of this circuit is $O(N\log N)$.
Figure 2: Quantum circuit illustrating computation of the determinant of a $4\times 4$ matrix. We omit superscript $(D)$ at $|\Phi_{k}\rangle$, $k=1,\dots,4$, and at $|\Psi_{out}\rangle$ for brevity. Operators $H_{A}$ and $X_{S}$ can be applied in parallel since they effect on different qubits.
III Inverse matrix
It is remarkable that the minor modification of the algorithm for calculating the determinant of a square $N\times N$ matrix in Sec.II allows to calculate the inverse of a square $(N-1)\times(N-1)$ matrix. Moreover, this modification leads to a minor increase in the depth of the algorithm which becomes $O(N^{2}\log N)$ instead of $O(N\log^{2}N)$ in the algorithm for calculating the determinant.
III.1 Preliminary arguments
We impose a special structure on the elements in the first row and first column of the matrix $\tilde{A}$ given in (6):
$a_{i0}=q,\;\;i=1,\dots,N-1,\;\;a_{0j}=1/\sqrt{N},\;\;j=0,\dots,N-1,$ (23)
i.e.,
$\tilde{A}=\left(\begin{array}[]{cccc}\frac{1}{\sqrt{N}}&\frac{1}{\sqrt{N}}&\cdots&\frac{1}{\sqrt{N}}\cr q&a_{11}&\cdots&a_{1(N-1)}\cr q&a_{21}&\cdots&a_{2(N-1)}\cr\cdots&\cdots&\cdots&\cdots\cr q&a_{(N-1)1}&\cdots&a_{(N-1)(N-1)}\end{array}\right).$ (29)
The algorithm discussed below allows to invert the $(N-1)\times(N-1)$ matrix $A$,
$A=\left(\begin{array}[]{ccc}a_{11}&\cdots&a_{1(N-1)}\cr a_{21}&\cdots&a_{2(N-1)}\cr\cdots&\cdots&\cdots\cr a_{(N-1)1}&\cdots&a_{(N-1)(N-1)}\end{array}\right),$ (34)
(rather than the matrix $\tilde{A}$ itself) which is embedded into $\tilde{A}$.
According to the standard definition, the element $A^{-1}_{ji}$ of the inverse matrix is expressed in terms of the minor $M_{ij}$ obtained by removing the $i$th row and $j$th column of $A$:
$A^{-1}_{ji}=\frac{(-1)^{i+j}M_{ij}}{\det(A)}.$ (35)
From another point of view, the minor $M_{ij}$ appears among the terms included in $\det(\tilde{A})$. In fact, one of the terms in this determinant is following (remove the $(i+1)$th row and 1st column from $\tilde{A}$; notice that $(i+1)$th row contains the elements $q$, $a_{i1}$, $a_{i3}$, …, $a_{i(N-1)}$):
$\frac{(-1)^{i+2}q}{\sqrt{N}}\left|\begin{array}[]{ccc}1&\cdots&1\cr a_{11}&\cdots&a_{1(N-1)}\cr\cdots&\cdots&\cdots\cr a_{(i-1)1}&\cdots&a_{(i-1)(N-1)}\cr a_{(i+1)1}&\cdots&a_{(i+1)(N-1)}\cr\cdots&\cdots&\cdots\cr a_{(N-1)1}&\cdots&a_{(N-1)(N-1)}\end{array}\right|,$ (43)
where we take into account that $A_{ij}=a_{ij}$, $i,j=1,\dots,N-1$. In turn, this term includes the following term (remove the 1st row and $j$th column from the determinant (43); the $j$th column contains the elements $1$, $a_{1j}$, $a_{2j}$, …, $a_{(N-1)j}$):
$-\frac{(-1)^{i+j}q}{\sqrt{N}}\left|\begin{array}[]{cccccc}a_{11}&\cdots&a_{1(j-1)}&a_{1(j+1)}&\cdots&a_{1(N-1)}\cr\cdots&\cdots&\cdots&\cdots&\cdots&\cdots\cr a_{(i-1)1}&\cdots&a_{(i-1)(j-1)}&a_{(i-1)(j+1)}&\cdots&a_{(i-1)(N-1)}\cr a_{(i+1)1}&\cdots&a_{(i+1)(j-1)}&a_{(i+1)(j+1)}&\cdots&a_{(i+1)(N-1)}\cr\cdots&\cdots&\cdots\cr a_{(N-1)1}&\cdots&a_{(N-1)(j-1)}&a_{(N-1)(j+1)}&\cdots&a_{(N-1)(N-1)}\end{array}\right|=$ (50)
$-\frac{q}{\sqrt{N}}(-1)^{i+j}M_{ij}=-\frac{\det(A)q}{\sqrt{N}}A^{-1}_{ji}.$
Thus, each element $A^{-1}_{ji}$ of the inverse matrix is included into $\det(\tilde{A})$ up to a factor which is the same for all elements. All what remains is to extract all the terms composing the element $A^{-1}_{ji}$ from those composing $\det(\tilde{A})$ and encode them into the proper state of some quantum subsystem. We show that this can be simply realized.
III.2 Basic algorithm
The circuit for the algorithm is presented in Fig.3.
Figure 3: Structure of a quantum circuit for matrix inversion. Ancillae are green on line.
According to (7), the pure state $|\Phi_{0}\rangle$ of the whole system is given in (9) with normalization (8) and constraints (23) for the elements in the first row and first column of the matrix $\tilde{A}$, see (29). From this state, we select only those terms that compose the determinant of $\tilde{A}$ and therefore contain the terms that appear in the resulting inverse matrix $A^{-1}$:
$|\Phi_{0}\rangle=\sum_{{k_{0},\dots,k_{N-1}=0}\atop{k_{0}\neq\dots\neq k_{N-1}}}^{N-1}a_{0k_{0}}\dots a_{(N-1)k_{N-1}}|k_{0}\rangle_{S_{0}}\cdots|k_{N-1}\rangle_{S_{N-1}}+|g_{1}\rangle_{S}$ (51)
while the garbage $|g_{1}\rangle_{S}$ will be removed at the last step of the algorithm after ancilla measurement. However, not all terms in the first part of $|\Phi_{0}\rangle$ are needed in our algorithm. We have to keep only those of them that include the 1st power of the factor $q$, (see constraints (23), i.e., exactly one state $|0\rangle_{S_{j}}$, $11\leq j\leq N-$ 1, can appear in the product $|k_{1}\rangle_{S_{1}}\dots|k_{N-1}\rangle_{S_{N-1}}$). We also have to associate each of the probability amplitudes in (51) with the appropriate element $A^{-1}_{ij}$ of the inverse matrix. To do so, we introduce two $n$-qubit subsystems $R$ and $C$ whose basis states serve to enumerate, respectively, rows and columns of the inverse matrix $A^{-1}$.
To select the terms that compose $A^{-1}_{ji}$ from the superposition state we use the projector
$P_{ij}=|0\rangle_{S_{i}}\;{}_{S_{i}}\langle 0|\otimes|j\rangle_{S_{0}}\;{}_{S_{0}}\langle j|,\;\;i,j=1,\dots,N-1.$ (52)
The idea is that this projector allows labeling the 1st and $(j+1)$th ($j=0,\dots,N-1$) columns and the 1st and $(i+1)$th ($10\leq i\leq N-$ 1) rows in the matrix $\tilde{A}$ (29). As the result, we select those terms from the superposition state whose probability amplitude compose $M_{ij}$.
Now we can construct the controlled operator
$W_{SRC}^{(ij)}=P_{ij}\otimes\sigma^{(x)}_{R^{(j)}C^{(i)}}+(I_{S}-P_{ij})\otimes I_{RC}.$ (53)
This operator conditionally applies $\sigma^{(x)}_{R^{(j)}C^{(i)}}$ which is the product of $\sigma^{(x)}$ (acting on the subsystems $R$ and $C$) such that $\sigma^{(x)}_{R^{(j)}C^{(i)}}|0\rangle_{R}|0\rangle_{C}=|j\rangle_{R}|i\rangle_{C}$. The depth of $W_{SRC}^{(ij)}$ is $O(n)$. To establish this estimation we use the fact that the controlled operator $W_{SRC}^{(ij)}$ with $n$-qubit controlling register can be represented in terms of Toffoli gates by a circuit of depth $O(n)$ KShV and require that all $\sigma^{(x)}$ operators in $\sigma^{(x)}_{R^{(j)}C^{(i)}}$ can be applied simultaneously. Applying the operator ${\hat{W}}_{SRC}=\prod_{i=1}^{N-1}\prod_{j=1}^{N-1}W_{SRC}^{(ij)}$ (whose depth is $O(N^{2}n)$) to $|\Phi_{0}\rangle|0\rangle_{R}|0\rangle_{C}$ we obtain
$|\Phi_{1}\rangle=\hat{W}_{SRC}|\Phi_{2}\rangle|0\rangle_{R}|0\rangle_{C}=$ (54)
$\frac{1}{\sqrt{N}}\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}\sum_{{0\neq k_{1}\neq\dots\neq k_{i-1}}\atop{\neq k_{i+1}\neq\dots\neq k_{N-1}\neq j}}a_{1k_{1}}\dots a_{(i-1)k_{i-1}}a_{i0}a_{(i+1)k_{i+1}}\dots a_{(N-1)k_{N-1}}\times$
$|j\rangle_{S_{0}}|k_{1}\rangle_{S_{1}}\dots|k_{i-1}\rangle_{S_{i-1}}|0\rangle_{S_{i}}|k_{i+1}\rangle_{S_{i+1}}\dots|k_{N-1}\rangle_{S_{N-1}}|j\rangle_{R}|i\rangle_{C}+|g_{2}\rangle_{SRC},$
where $a_{i0}=q$. Now we have to arrange the proper sign of each term in the first part of (54). According to Sec.III.1, this can be done via calculating the determinant of $\tilde{A}$ using the algorithm proposed in Sec.II. Applying the transformation $W^{(1)}_{SA}$ given in (15) to $|\Phi_{1}\rangle$ we obtain.
$|\Phi_{2}\rangle=W^{(1)}_{SA}|\Phi_{1}\rangle|0\rangle_{A}$ (55)
$=-\frac{q\det(A)}{2^{(\tilde{N}+n)/2}}\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}A^{-1}_{ji}|j\rangle_{R}|i\rangle_{C}|0\rangle_{S_{0}}|1\rangle_{S_{1}}\otimes|{N-1}\rangle_{S_{N-1}}|0\rangle_{A}+|g_{1}^{(D)}\rangle_{SA}.$
Next, applying the operator $X_{S}=\prod_{j=1}^{N-1}X_{S_{j}}(j)$ given in (18) to $|\Phi_{2}\rangle$ we obtain
$|\Phi_{3}\rangle=X_{S}|\Phi_{2}\rangle$ (56)
$=-\frac{q\det(A)}{2^{(\tilde{N}+n)/2}}\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}A^{-1}_{ji}|j\rangle_{R}|i\rangle_{C}|0\rangle_{S}|0\rangle_{A}+|g_{2}^{(D)}\rangle_{SA},$
Finally, applying $W^{(2)}_{SAB}$ (20) to $|\Phi_{3}\rangle|0\rangle_{B}$ we obtain
$|\Phi_{4}\rangle=W^{(2)}_{SAB}|\Phi_{3}\rangle|0\rangle_{B}$ $=$ $\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}a_{ji}|j\rangle_{R}|i\rangle_{C}|0\rangle_{S}|0\rangle_{A}|1\rangle_{B}+|g\rangle_{SA}|0\rangle_{B},$ (57)
$a_{ji}$ $=$ $-\frac{q\det(A)A^{-1}_{ji}}{2^{(\tilde{N}+n)/2}},$ (58)
The depth of the operator $D_{SAB}=W^{(2)}_{SAB}X_{S}W^{(1)}_{SA}$ (under certain assumptions on the physical realization of the operator $W_{k}$, see Eq.(13)) is $O(Nn^{2})<O(N^{2}n)$ (the depth of $\hat{W}_{SRC}$). After measuring the ancilla $B$ to get the output $|1\rangle$ with the probability $\frac{q^{2}|\det(A)|^{2}}{2^{\tilde{N}+n}}G^{2}\sim 2^{-O(\tilde{N})}=2^{-O(Nn)})$, $G=\sqrt{\sum_{i,j=1}^{N-1}|A^{-1}_{ij}|^{2}}$, we get the desired result
$|\Phi_{5}\rangle=|\Psi_{out}\rangle|0\rangle_{S}|0\rangle_{A}.$ (59)
where
$|\Psi_{out}\rangle$ $=$ $G^{-1}\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}A^{-1}_{ji}|j\rangle_{R}|i\rangle_{C}=G^{-1}|A^{-1}\rangle.$ (60)
The state $|\Psi_{out}\rangle$ encodes $A^{-1}$ in the normalized form, i.e., $|\Psi_{out}\rangle\langle\Psi_{out}|=1$, while the state $|A^{-1}\rangle=\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}A^{-1}_{ji}|j\rangle_{R}|i\rangle_{C}$ is not normalized. The depth of discussed algorithm for matrix inversion is determined by the operator $\hat{W}_{SRC}$ and equals $O(N^{2}\log N)$. The circuit includes $N\log N$ qubits of the subsystem $S$, $\tilde{N}=O(N\log N)$ qubits of the ancilla $A$, one qubit of the ancilla $B$ and $2\log N$ qubits of $R$ and $C$, thus the size of this circuit approximately equals to the size of the circuit for calculating the determinant which is $O(N\log N)$.
IV Application: solving systems of linear algebraic equations
Having constructed the inverse matrix $A^{-1}$ encoded into the state of the subsystem $R\cup C$ we obtain a tool for solving a system of linear algebraic equations represented in the following matrix form:
$A\bm{x}=\bm{b}.$ (61)
Its solution reads
$\bm{x}=A^{-1}\bm{b}.$ (62)
Therefore, we have to apply the multiplication algorithm proposed in Ref.ZQKW_arxiv2023 (see also Appendix, Sec.VI.1, for its modified version) to the matrices $A^{-1}$ and $\bm{b}$. The structure of the circuit is shown in Fig.4. More detailed circuit is presented in Fig.5.
Figure 4: Structure of the circuit solving a system of linear algebraic equations. Both blocks share the same ancillae, which are not shown in figure.
Notice that we can use the same ancillae in both blocks of this circuit (ancillae are not shown in Fig.4). Moreover, the subsystem $S$ encoding the matrix $A$ can be used as ancilla in further calculations.
In this case we do not measure ancillae $B$ (unlike Sec.III.2) and thus start with the state $|\Phi_{4}\rangle$ given in (57). The vector $\bm{b}$ is encoded as
$|b\rangle=\sum_{j=0}^{N-1}b_{j}|j\rangle_{b},\;\;b_{0}=0,\;\;\sum_{j=1}^{N-1}|b_{j}|^{2}=1.$ (63)
The zero value for $b_{0}$ is prescribed by the encoding of $A^{-1}$ in (60), where the 00-states of $R$ and $C$ are not used. Out of system $S\cup A$ (which form ancilla in Fig.5), we take the one-qubit ancilla $\tilde{B}$ needed in the multiplication algorithm and neglect other qubits. We also apply the operator $X_{B}=\sigma^{(x)}$ to invert the state of $B$. Thus, the state of the whole system reads
$|\Phi^{(L)}_{5}\rangle$ $=$ $X_{B}|\Phi_{4}\rangle|b\rangle|0\rangle_{\tilde{B}}=a_{jl}b_{i}|j\rangle_{R}|l\rangle_{C}|i\rangle_{b}|0\rangle_{B}|0\rangle_{\tilde{B}}+|g_{1}\rangle_{RCb}|1\rangle_{B}|0\rangle_{\tilde{B}}.$ (64)
Now we refer to the Appendix, Sec.VI.1, where the matrix multiplication algorithm is described. Applying the result of this Appendix we identify $K=N$, $M=1$ and replace the subsystems $R_{1}$, $C_{1}$, $R_{2}$ with $R$, $C$, $b$, while the subsystem $C_{2}$ is not used because the second matrix in our multiplication is just a column. Thus, $W^{(1)}_{C_{1}R_{2}}$ in (71) becomes $W^{(1)}_{Cb}$, $W^{(2)}_{C_{1}}$ in (73) becomes $W^{(2)}_{C}$, while the operator $W^{(3)}_{C_{1}R_{2}\tilde{B}}$ in (74) must be modified including the projector acting on $B$, i.e. the new operator reads:
$W^{(3)}_{CBb\tilde{B}}=P_{CBb}\otimes\sigma^{(x)}_{\tilde{B}}+(I_{CBb}-P_{CBb})\otimes I_{\tilde{B}},\;\;P_{CBb}=|0\rangle_{C}|0\rangle_{B}|0\rangle_{b}\;{{}_{C}}\langle 0|_{B}\langle 0|{{}_{b}}\langle 0|.$ (65)
Applying the operator $W^{(3)}_{CBb\tilde{B}}W^{(2)}_{C}W^{(1)}_{Cb}$ to $|\Phi^{(L)}_{5}\rangle$ (64) we obtain
$|\Phi^{(L)}_{6}\rangle=W^{(3)}_{CBb\tilde{B}}W^{(2)}_{C}W^{(1)}_{Cb}|\Phi^{(L)}_{5}\rangle=$ (66)
${\frac{1}{2^{n/2}}}\left({\sum_{j_{1}=1}^{N-1}}\sum_{j=1}^{N-1}a_{j_{1}j}b_{j}|j_{1}\rangle_{R}|0\rangle_{C}|0\rangle_{b}\right)|0\rangle_{B}|1\rangle_{\tilde{B}}+$
$|g_{2}\rangle_{RCb}|1\rangle_{B}|0\rangle_{\tilde{B}}.$
Performing measurement over $\tilde{B}$ with the output $|1\rangle_{\tilde{B}}$ we remove garbage and obtain
$|\Phi^{(L)}_{7}\rangle=|\Psi^{(L)}_{out}\rangle\,|0\rangle_{C}|0\rangle_{b}|0\rangle_{B},$
$|\Psi^{(L)}_{out}\rangle=G^{{-1}}{\sum_{j_{1}=1}^{N-1}}\sum_{j=1}^{N-1}a_{j_{1}j}b_{j}|j_{1}\rangle_{R}=G^{{-1}}|x\rangle_{R},$
where $|x\rangle_{R}=\sum_{j=1}^{N-1}a_{j_{1}j}b_{j}|j_{1}\rangle_{R}$ is an unnormalized vector, the normalization $G{=(\sum_{l,j}|a_{lj}b_{j}|^{2})^{1/2}}$ is known from the probability of the above measurement which equals $G^{2}/2^{n}$, where $G^{2}\sim\frac{1}{2^{\tilde{N}}}\sim\frac{1}{2^{O(Nn)}}$ due to the expression (58) for $a_{jl}$. Thus, the probability is $\sim 1/2^{\tilde{N}+n}=2^{-O(Nn)}$. The result is stored in the register $R$. The depth of the whole algorithm is defined by the depth of the subroutine inverting $A$ and thus equals $O(N^{2}n)$. In comparison with the circuit for inverse matrix, the circuit in Fig. 5 includes $\log N$ qubits encoding the column $\bm{b}$ and thus remains essentially $O(N\log N)$.
Figure 5: Circuit for the algorithm solving a system of linear algebraic equations (61). Circuit of matrix inversion (up to the operator $W^{(2)}_{SAB}$) is denoted by a block $A^{-1}$. All input qubits of the matrix-inversion algorithm (system $S$) become ancillae qubits. The output of the matrix-inversion algorithm (systems $R$ and $C$) together with system $b$ encoding the vector $\bm{b}$ become the input of the matrix-multiplication block. The output of the whole algorithm is encoded in the state $|x\rangle_{R}$ of the system $R$. All ancillae qubits are green on line. We omit the superscript $(L)$ at $|\Phi_{k}\rangle$, $k=5,6,7$ and at $|\Psi_{out}\rangle$ for brevity.
V Conclusions
We propose two quantum algorithms for matrix manipulation, which are calculating the determinant and matrix inversion. Both are based on special row-wise encoding of the matrix elements in the pure states of quantum subsystems.
The algorithm for computing the determinant of a square matrix is discussed in details. It uses multiqubit control-SWAP and control-Z operations. The depth of this algorithm is $O(N\log^{2}N)$. This algorithm utilized the subroutine realizing the completely antisymmetric tensor $\varepsilon$ (operator $W^{(1)}_{SA}$) that may have relevance in other algorithms. The outcome of this algorithm is split into two components: the absolute value of the determinant of the matrix $A$, denoted as $|\det(A)|$, and the argument of the determinant of the matrix $A$, denoted as $\arg(\det(A))$. The first part is obtained from the measurement of an ancilla, (although the probability of the needed outcome in such measurement decreases exponentially with $N$), while the argument of the determinant remains embedded in the probability amplitude of the output state $|\Psi^{(D)}_{out}\rangle$. The size of the subsystem $S$ encoding the matrix $A$ is $N\log N$, the size of ancillae $A_{k}$, $k=0,\dots,N-2$, is $\tilde{N}=O(N\log N)$. Both these sizes serve to estimate the size of the whole algorithm which is $O(N\log N)$.
It is remarkable that the $(N-1)\times(N-1)$-matrix inversion algorithm is essentially algorithm for calculating the determinant of the $N\times N$ matrix obtained by replacing the first row by the row of $1/\sqrt{N}$ and the first column by the column of $q$’s (up to the diagonal element). This special encoding provides the superposition of all products of matrix elements which are necessary for calculating minors included into the definition of the inverse matrix. The algorithm distributes these products with proper signs among the appropriate inverse-matrix elements which are encoded into the quantum state of certain subsystem thus forming the output state. The depth of the algorithm is $O(N^{2}\log N)$.
As an application of inversion we develop the algorithm for solving systems of linear algebraic equations. This algorithm consists of two blocks: matrix inversion and matrix multiplication. At that, the matrix inversion is most resource-consuming block which finally defines the depth of the algorithm (which is $O(N^{2}\log N)$), and probability of required ancilla state after measurement, $\sim 2^{-O(N\log N)}$. We have to notice that the depths of the algorithms for calculating the determinant and matrix inverse were obtained under assumption that the $U$-, $V$-operators in (13) can be applied in parallel. Otherwise we have to change the depths $O(N\log^{2}N)$ and $O(N^{2}\log\,N)$ to $O(N^{2}\log^{2}N)$ in both cases.
We emphasize that the matrix-multiplication algorithm is modified in comparison with that proposed in ZQKW_arxiv2023 . The detailed description of the modified version is presented in Appendix, Sec. VI.1.
One notable aspect is the use of ancillae to gather all the garbage generated during the calculation, which can then be discarded with a measurement of a one-qubit ancilla. The weak point of both algorithms is probability of obtaining the required state of ancilla after measuring. This probability is $\sim 2^{-O(N\log N)}$ in all three algorithms.
Combining the algorithm of matrix inversion with the algorithm of matrix multiplication we demonstrate the usage of two types of matrix encoding into the state of a quantum system. First one is the row-wise encoding used in the input of $N\times N$ matrix inversion algorithm (for simplicity, we assume that $N=2^{n}$). It requires $N$ subsystems of $\log N$ qubits each, and the state of $i$th subsystem encodes the elements of the $i$th row of the matrix. The result of matrix inversion is stored in the two subsystems of $\log N$ qubits each. These subsystems enumerate rows and columns respectively. This form of encoding of the inverse matrix allows to perform its multiplication with the column $\bm{b}$ properly encoded in the state of $\log N$-qubit subsystem to obtain the solution $\bm{x}$ of the linear system. This solution is a column whose elements are encoded into the state of a $\log N$-qubit subsystem $R$.
Acknowledgements. This project is supported by the National Natural Science Foundation of China (Grants No. 12031004, No. 12271474, and No. 61877054), and the Fundamental Research Foundation for the Central Universities (Project No. K20210337) and Zhejiang University Global Partnership Fund, 188170+194452119/003. The work was partially funded by a state task of Russian Fundamental Investigations (State Registration No. 124013000760-0).
VI Appendix
VI.1 Matrix multiplication
We present an algorithm for multiplying an $N\times K$ matrix $A^{(1)}$ by a $K\times M$ matrix $A^{(2)}$ with the elements $A^{(i)}=\{a^{(i)}_{jk}\}$, $i=1,2$, assuming $N=2^{n}$, $K=2^{k}$, $M=2^{m}$ with positive integers $n$, $k$, $m$. The algorithm below is an alternative to that proposed in ZQKW_arxiv2023 . Both the depth and size of the algorithm are less than those parameters in ZQKW_arxiv2023 , although the estimations for them remains the same: depth is $O(\log K)$ and size is $O(\log K+\log N+\log M)$.
We first introduce one register of $n$ qubits, two registers of $k$ qubits and one register of $m$ qubits which enumerate rows and columns of both matrices. The pure states encoding the elements of matrices are
$|\Psi_{i}\rangle$ $=$ $\sum_{jl}a^{(i)}_{jl}|j\rangle_{R_{i}}|l\rangle_{C_{i}},\;\;\;i=1,2,$ (67)
with normalization conditions
$\sum_{jl}|a^{(i)}_{jl}|^{2}=1,\;\;\;i=1,2.$ (68)
The initial state of the whole system reads
$|\Phi_{0}^{(m)}\rangle=|\Psi_{1}\rangle\otimes|\Psi_{2}\rangle=$ (69)
$\sum_{j_{1}=0}^{N-1}\sum_{l_{1},j_{2}=0}^{K-1}\sum_{l_{2}=0}^{M-1}a^{(1)}_{j_{1}l_{1}}a^{(2)}_{j_{2}l_{2}}|j_{1}\rangle_{R_{1}}|l_{1}\rangle_{C_{1}}|j_{2}\rangle_{R_{2}}|l_{2}\rangle_{C_{2}}.$
From the superposition state $|\Phi_{0}\rangle$, we have to select terms with $l_{1}=j_{2}$. For that purpose we apply the CNOTs $W_{j}$ to the $j$th qubits of $C_{1}$ and $R_{2}$,
$W_{j}$ $=$ $|1\rangle_{C_{1},j}\;{}_{C_{1},j}\langle 1|\otimes\sigma^{(x)}_{R_{2},j}+|0\rangle_{C_{1},j}\;{}_{C_{1},j}\langle 0|\otimes I_{R_{2},j}.$ (70)
where $I_{R_{2},j}$ is the identity operators acting on the $j$th qubit of the subsystem $R_{2}$. We have chosen the qubits of $C_{1}$ as control ones. However, if $C_{1}$ and $R_{2}$ are exchanged here and below in the description of the algorithm, it remains correct. All operators $W_{j}$, $j=1,\dots,k$ commute with each other. Applying the operator $W^{(1)}_{C_{1}R_{2}}$,
$W^{(1)}_{C_{1}R_{2}}=\prod_{j=1}^{k}W_{j},$ (71)
to $|\Phi_{0}^{(m)}\rangle$ we obtain:
$|\Phi_{1}^{(m)}\rangle=W^{(1)}_{C_{1}R_{2}}|\Phi_{0}^{(m)}\rangle=$ (72)
${\sum_{j_{1}=0}^{N-1}}\sum_{j=0}^{K-1}\sum_{l_{2}=0}^{M-1}a^{(1)}_{j_{1}j}a^{(2)}_{jl_{2}}|j_{1}\rangle_{R_{1}}|j\rangle_{C_{1}}|0\rangle_{R_{2}}|l_{2}\rangle_{C_{2}}+$
$|g_{1}^{(m)}\rangle_{R_{1}C_{1}R_{2}C_{2}},$
selecting terms with $|0\rangle_{R_{2}}$ from the superposition state $|\Phi_{1}^{(m)}\rangle$ (the other terms are by definition garbage; their sum is denoted by $|g_{1}\rangle$ above). Now we apply the Hadamard transformations
$W^{(2)}_{C_{1}}=H^{\otimes k}$ (73)
to $C_{1}$ and select the terms with $|0\rangle_{C_{1}}|0\rangle_{R_{2}}$:
$|\Phi_{2}^{(m)}\rangle=W^{(2)}_{C_{1}}|\Phi_{1}^{(m)}\rangle=$
$\frac{1}{2^{k/2}}{\sum_{j_{1}=0}^{N-1}}\sum_{j=0}^{K-1}\sum_{l_{2}=0}^{M-1}a^{(1)}_{j_{1}j}a^{(2)}_{jl_{2}}|j_{1}\rangle_{R_{1}}|0\rangle_{C_{1}}|0\rangle_{R_{2}}|l_{2}\rangle_{C_{2}}$
$+|g_{2}^{(m)}\rangle_{R_{1}C_{1}R_{2}C_{2}}.$
Here the first term contains the desired matrix product. Next, to label the new garbage, we prepare a one-qubit ancilla $\tilde{B}$ in the ground state $|0\rangle_{\tilde{B}}$, introduce the projector $P_{C_{1}R_{2}}=|0\rangle_{C_{1}}|0\rangle_{R_{2}}\;{{}_{C_{1}}}\langle 0|{{}_{R_{2}}}\langle 0|$ and the controlled operator $W^{(3)}_{C_{1}R_{2}\tilde{B}}$,
$W^{(3)}_{C_{1}R_{2}\tilde{B}}=P_{C_{1}R_{2}}\otimes\sigma^{(x)}_{\tilde{B}}+(I_{C_{1}R_{2}}-P_{C_{1}R_{2}})\otimes I_{\tilde{B}},$ (74)
of the depth $O(k)$ with $({2}k)$-qubit control register. Applying this operator to $|\Phi_{2}^{(m)}\rangle|0\rangle_{\tilde{B}}$ we obtain
$|\Phi_{3}^{(m)}\rangle=W^{(3)}_{C_{1}R_{2}\tilde{B}}|\Phi_{2}^{(m)}\rangle|0\rangle_{\tilde{B}}=$ (75)
${\frac{1}{2^{k/2}}}\left({\sum_{j_{1}=0}^{N-1}}\sum_{j=0}^{K-1}\sum_{l_{2}=0}^{M-1}a^{(1)}_{j_{1}j}a^{(2)}_{jl_{2}}|j_{1}\rangle_{R_{1}}|0\rangle_{C_{1}}|0\rangle_{R_{2}}|l_{2}\rangle_{C_{2}}\right)|1\rangle_{\tilde{B}}+$
$|g_{2}\rangle_{R_{1}C_{1}R_{2}C_{2}}|0\rangle_{\tilde{B}}.$
Performing measurement over $\tilde{B}$ with the output $|1\rangle_{\tilde{B}}$, we remove the garbage and obtain
$|\Phi_{4}^{(m)}\rangle=|\Psi_{out}^{(m)}\rangle\,|0\rangle_{C_{1}}|0\rangle_{R_{2}},$
$|\Psi_{out}^{(m)}\rangle=G^{{-1}}{\sum_{j_{1}=0}^{N-1}}\sum_{j=0}^{K-1}\sum_{l_{2}=0}^{M-1}a^{(1)}_{j_{1}j}a^{(2)}_{jl_{2}}|j_{1}\rangle_{R_{1}}|l_{2}\rangle_{C_{2}},$
where the normalization factor $G{=(\sum_{j_{1},j,l_{2}}|a^{(1)}_{j_{1}j}a^{(2)}_{jl_{2}}|^{2})^{1/2}}$ is known from the probability of the above measurement which equals $G^{2}/2^{k}\sim O(2^{-k})$. The result of the multiplication is stored in the state of the subsystem $R_{1}\cup C_{2}$. From the above analysis we conclude that the depth of the whole protocol is defined by the operator $W^{(3)}_{C_{1}R_{2}\tilde{B}}$ and equals $O(k)=O(\log(K))$. The circuit is given in Fig. 6, its size is $12k+n+m+$ 1.
Figure 6: (a) Circuit for the matrix multiplication algorithm; (b) Notation for multiqubit CNOT. We omit superscript $(m)$ at $|\Phi_{k}\rangle$, $k=0,\dots,4$, and at $|\Psi_{out}\rangle$ for brevity.

### v4

@@
-Quantum algorithms for calculating determinant and inverse of matrix and solving linear algebraic systems
+Purely quantum algorithms for calculating determinant and inverse of matrix and solving linear algebraic systems
 Abstract
-We propose quantum algorithms for calculating the determinant and inverse of an $(N-1)\times(N-1)$ matrix (depth is $O(N^{2}\log N)$) which is a simple modification of the algorithm for calculating the determinant of an $N\times N$ matrix (depth is $O(N\log^{2}N)$.
-The basic idea is to encode each row of the matrix into a pure state of some quantum system.
+We propose quantum algorithms that are fundamentally quantum in nature for the computation of the determinant and inverse of an $(N-1)\times(N-1)$ matrix, with a computational depth of $O(N^{2}\log N)$.
+This approach is a straightforward modification of the existing algorithm for determining the determinant of an $N\times N$ matrix, which operates with a depth of $O(N\log^{2}N)$.
+The core concept involves encoding each row of the matrix into a pure state of a quantum system.
 In addition, we use the representation of the elements of the inverse matrix in terms of algebraic complements.
-Combination of this algorithm with the quantum algorithm for matrix multiplication, proposed in our previous work, yields the algorithm for solving systems of linear algebraic equations (depth is $O(N\log^{2}N)$.
-Measurement of the ancilla state with output 1 (probability is $\sim 2^{-O(N\log N)}$) removes the garbage acquired during calculation.
-Appropriate circuits for all three algorithm are presented and have the same estimation $O(N\log N)$ for the size.
-Keywords: quantum algorithm, quantum circuit, determinant, inverse matrix, linear system, depth of algorithm, quantum measurement
+This algorithm, in conjunction with the matrix multiplication algorithm discussed in the appendix, facilitates solving systems of linear algebraic equations with a depth of $O(N\log^{2}N)$.
+The measurement of the ancilla state yielding an output of 1 occurs with a probability of approximately $2^{-O(N\log N)}$, and eliminates the extraneous data generated during the computation.
+We provide suitable circuit designs for all three algorithms, each estimated to require $O(N\log N)$ in terms of space, specifically the number of qubits utilized in the circuit.
+Keywords: quantum algorithm, quantum circuit, determinant and inverse of matrix, linear algebraic systems, depth of algorithm, quantum measurement
 I Introduction
-Quantum algorithms that exploit principles of quantum physics can be run on quantum computers and promise to offer considerable advantages over classical counterparts.
+Quantum algorithms exploiting principles of quantum physics can be run on quantum computers and promise to offer considerable advantages over classical counterparts.
 Algorithms for quantum calculations are an increasingly prominent field in quantum information research.
-Notable examples include Shor’s factoring algorithm Sho1 ; Sho2 , Deutch parallelism Deu , Grover’s search algorithm Gro , Quantum Fourier transform QFT1 ; QFT2 ; QFT3 and Quantum phase estimation Wang ; QFT3 .
+Notable examples include Shor’s factoring algorithm Sho1 ; Sho2 , Deutsch parallelism Deu , Grover’s search algorithm Gro , Quantum Fourier transform QFT1 ; QFT2 ; QFT3 and Quantum phase estimation Wang ; QFT3 .
 Further advancements include the HHL algorithm for solving systems of linear equations HHL ; HHL1 ; HHL2 ; HHL3 ; HHL4 ; HHL5 ; HHL6 ; HHL7 , in which the crucial point is the conditional rotation of the state of the one-qubit ancilla.
@@
 HHL algorithm exponentiates a Hermitian matrix of the considered linear system, which can be performed via the Trotterization formula T ; AT .
-There is extensive literature devoted to improvement of this algorithm, especially its scale and complexity BHMT ; CKS ; GSLW ; LC ; LYC ; AL ; JRS ; SSO ; AlL ; TAWL .
+There is extensive literature devoted to improvement of this algorithm, especially its scaling and complexity BHMT ; CKS ; GSLW ; LC ; LYC ; AL ; JRS ; SSO ; AlL ; TAWL .
 However, the principal problem of eigenvalue inversion is not resolved via quantum methods yet.
-For this purpose, for instance, the quantum state tomography can be involved to obtain the eigenvalues of the exponent of Hermitian operator (the matrix of the linear system) T .
-Nevertheless, the HHL algorithm combining the phase estimation and classical inversion of eigenvalues is very popular in realization of various algorithms of quantum computation and it is widely applicable in quantum machine learning A ; HHL2 ; RML .
+For this purpose, for instance, the quantum state tomography can be used to obtain the eigenvalues of the exponent of Hermitian operator (the matrix of the linear system) T .
+Nevertheless, the HHL algorithm combining the phase estimation and classical inversion of eigenvalues is very popular in implementation of various algorithms of quantum computation and it is widely applicable in quantum machine learning A ; HHL2 ; RML .
 In particular, it is applied in the least-square linear-regression algorithms WBL ; SSP ; Wang2 working with large data sets, in quantum algorithms for singular-value decomposition GSLW ; G , in algorithms for solving linear B and nonlinear LKKLTC differential equations.
-We also have to refer to algorithms for matrix manipulations based on the Trotterization method and the Baker-Champbell-Housdorff formula for exponentiating matrices ZhaoL1 .
+We also refer to algorithms for matrix manipulations based on the Trotterization method and the Baker-Champbell-Hausdorff formula for exponentiating matrices ZhaoL1 .
 In GSLW , algebraic manipulations with matrices are implemented using unitary transformations that include the matrices as specific blocks (block-encoding model).
@@
 Later, in ZQKW_arxiv2023 , we proposed the implementation of unitary transformations corresponding to vector inner product, matrix sum, and product using multi-qubit conditional operators (reducible to Toffoli gates) and Hadamard transformations, including the ancilla measurement for garbage removal.
-The depth or complexity of these algorithms increases logarithmically with the matrix dimension, or even remains independent of it (in the case of matrix sum), demonstrating the advantage of these algorithms compared to their classical counterparts.
-In this paper, we propose the algorithm for matrix inversion with its application to solving systems of linear equations.
-The principal part of this algorithm is the algorithm for computing the determinant of a square matrix.
+The depth of these algorithms increases logarithmically with the matrix dimension, or even remains independent of it (in the case of matrix sum), demonstrating the advantage of these algorithms compared to their classical counterparts.
+In this paper, we propose quantum algorithms for determinant and inversion of matrix with their application to solving systems of linear equations.
+The principal part of these algorithms is the algorithm for computing the determinant of a square matrix.
 It is built on the ideas of Refs.
 QZKW_arxive2022 ; ZQKW_arxiv2023 :
-    Encode matrix elements into the states of a certain quantum subsystems through probability amplitudes in such a way that all necessary multiplications of matrix elements are performed automatically and appear in the probability amplitudes of the initial state of the whole system;
+    Encode matrix elements into the states of certain quantum subsystems through probability amplitudes in such a way that all necessary multiplications of matrix elements are performed automatically and appear in the probability amplitudes of the initial state of the whole system;
     Use multiqubit control-SWAP and control-Z operators together with ancilla to organize proper signs of each term in the superposition state and collect them in certain state of the considered system;
@@
 We follow the above steps to propose an algorithm for calculating the determinant of a square matrix whose $N$ rows are encoded into the pure states of $N$ separate quantum subsystems.
-Notice that creating such initial state is a special problem by itself, but we do not handle this problem here.
+Notice that creating such initial states is a special problem in itself, but we do not discuss this problem here.
 The state of the whole system results in all possible products of $N$ elements from different rows.
@@
 We propose such an algorithm and construct the quantum circuit realizing it.
-While developing the algorithm, our focus is on the sizes and depths of its specific blocks.
-Essentially, our algorithm is a quantum realization of the cofactor expansion classical algorithm, whose depth exponentially increases with matrix dimension.
-However, the resulting depth of our algorithm is $O(N\log^{2}N)$ (or $O(N^{2}\log N)$ in the worst case) with the size $O(N\log N)$.
-The final step of this algorithm is the one-qubit ancilla measurement, which removes all the garbage.
-The probability of obtaining the required result of the ancilla measurement (which is 1) exponentially decreases with the matrix dimension $N$, which requires exponentially increasing number of evaluations of the algorithm.
-This is a central problem that we plan to overcome in further work.
+While developing the algorithm, our focus is on the space and depths of its specific blocks.
+Essentially, our algorithm is a quantum realization of the cofactor expansion of the classical algorithm, whose depth exponentially increases with matrix dimension.
+However, the resulting depth of our algorithm is $O(N\log^{2}N)$ (or $O(N^{2}\log N)$ in the worst case) with space of $O(N\log N)$ qubits.
+The final step of this algorithm is the one-qubit ancilla measurement which removes all the garbage.
+The probability of obtaining the required result of the ancilla measurement (which is 1) diminishes exponentially as the matrix dimension $N$ increases, necessitating an exponentially larger number of algorithm evaluations.
+This is a central problem that we plan to investigate in future work.
 It is remarkable that the minor modification of the algorithm for calculating the determinant yields an alternative quantum algorithm for matrix inversion using its definition in terms of algebraic complements.
-We again use the row-wise encoding of the matrix elements into the state of a quantum system so that all necessary products of matrix elements appear automatically in the superposition state, thus one only needs to properly sort these products.
-The depth of this algorithm is $O(N^{2}\log N)$ (or $O(N^{2}\log^{2}N)$ in the worst case) and the size is $O(N\log N)$.
+We again use the row-wise encoding of the matrix elements into the state of a quantum system so that all necessary products of matrix elements appear automatically in the superposition state.
+One thus only needs to properly sort these products.
+The depth of this algorithm is $O(N^{2}\log N)$ (or $O(N^{2}\log^{2}N)$ in the worst case) and the space is $O(N\log N)$ qubits.
 Based on such inversion algorithm we develop an alternative algorithm for solving systems of linear algebraic equations.
 Thus, we do not use Trotterization methods and classical subroutine for inversion of eigenvalues.
-Therefore the proposed algorithm is completely quantum and it is an alternative to the inversion algorithm considered in HHL ; T ; TAWL .
-The estimations for the depth and size of this algorithm remain the same.
-In relation with the question of algorithm complexity, we shall mention that the special methods of classical manipulations with matrices allow to reduce the depth of the algorithms for determinant calculation to $O(\log^{2}N)$ Ber ; BAEM with the size $O(N^{3.496})$.
-In our case we use the Leibniz formula as a definition of the matrix determinant and definition of the inverse matrix in terms of the algebraic complements.
-Our quantum algorithms based on these definitions obviously reveal advantages in comparison with classical realizations of these definitions.
-It is likely that using specific classical algorithms for matrix manipulations and developing their quantum analogues lead to further improvement in the depth and size of quantum algorithms calculating the determinant and inverse matrix.
-This issue is not considered in our paper.
-The main novelty is that we are promoting the method for representation of the matrix elements as probability amplitudes of pure states and reveal profits out of such representation in application to the standard algorithms for matrix manipulations.
+Therefore, the proposed algorithm is purely quantum in nature and it gives an alternative avenue to the inversion algorithm considered in HHL ; T ; TAWL .
+The estimations for the depth and space of this algorithm remain the same.
+Regarding the question of algorithmic complexity, we remark that the specialized classical techniques of matrix manipulations facilitate a reduction in the depth of the algorithms for determinant calculation to $O(\log^{2}N)$ Ber ; BAEM with the size $O(N^{2.496})$.
+In our study, we employ the Leibniz formula as a definition for the matrix determinant and define the inverse matrix in terms of the algebraic complements.
+Our quantum algorithms based on these definitions demonstrate clear advantages over their classical counterparts.
+It is plausible that the adaptation of specific classical algorithms for matrix operations into their quantum equivalents could yield further enhancements in the efficiency and complexity of quantum algorithms designed for calculating determinants and inverse matrices.
+However, this particular aspect is not addressed in the present paper.
+The primary contribution of our work lies in promoting the representation of matrix elements as probability amplitudes of pure states, and we elucidate the benefits of this representation in relation to standard algorithms for matrix manipulations.
 The paper is organized as follows.
 The algorithm for calculating the matrix determinant is proposed in Sec.II.
-This algorithm is taken as a basis for the algorithm calculating the inverse matrix presented in Sec.
-III.
-Based on that results, the algorithm for solving systems of linear algebraic equations is proposed in Sec.
-IV.
+This algorithm is taken as a basis for the algorithm calculating the inverse matrix presented in Sec.III.
+Based on those results, the algorithm for solving systems of linear algebraic equations is proposed in Sec.IV.
 The main results are discussed in Sec.V.
@@
 $\tilde{A}=\left(\begin{array}[]{cccc}a_{00}&a_{01}&\cdots&a_{0(N-1)}\cr a_{10}&a_{11}&\cdots&a_{1(N-1)}\cr\cdots&\cdots&\cdots&\cdots\cr a_{(N-2)0}&a_{(N-2)1}&\cdots&a_{(N-2)(N-1)}\cr a_{(N-1)0}&a_{(N-1)1}&\cdots&a_{(N-1)(N-1)}\end{array}\right),$ (6)
-we encode each row of $N$ elements into the pure state $|\Psi_{j}\rangle$ of the $n$-qubit subsystem $S_{j}$,
+we encode each row consisting of $N$ elements into the pure state $|\Psi_{j}\rangle$ of the $n$-qubit subsystem $S_{j}$,
 $|\Psi_{j}\rangle=\sum_{k=0}^{N-1}a_{jk}|k\rangle_{S_{j}},\;\;j=0,1,\dots,N-1.$ (7)
-Representing rows of the matrix $\tilde{A}$ in the form of pure states (7) presumes $N$ obvious constraints for the matrix elements $a_{ij}$ predicted by the normalization of pure states:
+Representing rows of the matrix $\tilde{A}$ in the form of pure states (7) creates $N$ obvious constraints for the matrix elements $a_{ij}$ because of the normalization of pure states:
 $\sum_{k=0}^{N-1}|a_{jk}|^{2}=1,\;\;j=0,\dots,N-1.$ (8)
-These constraints can be smoothened by incorporating additional terms with different probability amplitudes into the states $|\Psi_{i}\rangle$, but this issue will be addressed in future study.
-The whole system is the composition of the subsystems, $S=\bigcup_{j}S_{j}$, and the pure state of the whole system reads
+These constraints can be circumvented by incorporating additional terms with different probability amplitudes into the states $|\Psi_{i}\rangle$, but this issue will be addressed in future study.
+The whole system is the union of the subsystems, $S=\bigcup_{j}S_{j}$, and the pure state of the whole system reads
 $|\Phi_{0}\rangle$ $=$ $|\Psi_{0}\rangle\cdots|\Psi_{N-1}\rangle$ (9)
-$=$ $\sum_{k_{0},\dots,k_{N-1}=0}^{N-1}a_{0k_{0}}\dots a_{(N-1)k_{N-1}}|k_{0}\rangle_{S_{0}}\cdots|k_{N-1}\rangle_{S_{N-1}},$
+$=$ $\sum_{k_{0},\dots,k_{N-1}=0}^{N-1}a_{0k_{0}}\dots a_{(N-1)k_{N-1}}$
+$\times$ $|k_{0}\rangle_{S_{0}}\cdots|k_{N-1}\rangle_{S_{N-1}},$
 which includes $N^{(S)}=nN$ qubits. In computing the determinant, we need only those terms in which all $k_{j}$ are different. We select these terms from the whole pure state of our system:
-$|\Phi_{0}^{(D)}\rangle$ $=$ $\sum_{k_{0}\neq\dots\neq k_{N-1}}a_{0k_{0}}\dots a_{(N-1)k_{N-1}}|k_{0}\rangle_{S_{0}}\cdots|k_{N-1}\rangle_{S_{N-1}}+|g_{1}\rangle,$ (10)
-where $|g_{1}\rangle$ collects all extra terms which form garbage to be removed later.
+$|\Phi_{0}\rangle$ $=$ $\sum_{{k_{0},\dots,k_{N-1}=0}\atop{k_{0}\neq\dots\neq k_{N-1}}}^{N-1}a_{0k_{0}}\dots a_{(N-1)k_{N-1}}$ (10)
+$\times$ $|k_{0}\rangle_{S_{0}}\cdots|k_{N-1}\rangle_{S_{N-1}}+|g_{1}\rangle_{S},$
+where $|g_{1}\rangle$ collects all extra terms which form the garbage to be removed later.
 We have to arrange the appropriate signs for each term of the first part in the right-hand side of (10).
 This can be achieved by setting the subsystem $S_{j}$ (in each term) into the state $|j\rangle$ using the SWAP operators together with sign-switching.
-This pair of operators performs permutation properly altering the sign of the term.
+This pair of operators performs permutation and changes the sign of the term.
 Thus, our purpose is to exchange the states of subsystems $S_{j}$ ($j=0,\dots,N-1$) so that the state of the subsystem $S_{k}$ becomes $|k\rangle_{S_{k}}$ in each term of the first part of superposition state (10).
@@
 The ancilla $A_{k}$ serves to save the index $j$ of the subsystem $S_{j}$ whose state is $|k\rangle$ with $j>k$.
-Therefore, the number of qubits $\tilde{n}_{k}$ in $A_{k}$ is less then $n=\log N$ unless $k=0$: $\tilde{n}_{k}\leq n$.
+Therefore, the number of qubits $\tilde{n}_{k}$ in $A_{k}$ is less than $n=\log N$ unless $k=0$: $\tilde{n}_{k}\leq n$.
 For instance, in Fig.
@@
 $V_{S_{j}A_{k}}$ $=$ $P_{S_{j}}^{(k)}\otimes Z_{k}(j)\sigma^{(x)}_{kj}+(I_{S_{j}}-P_{S_{j}}^{(k)})\otimes I_{A_{k}},$ (11)
-where $\sigma^{(x)}_{kj}$ includes the set of $\sigma^{(x)}$-operators applied to the ground state of the ancilla $A_{k}$ to form the state $|\tilde{j}\rangle_{A_{k}}$; $Z_{k}(j)$ changes the sign of the ancilla state $|\tilde{j}\rangle_{A_{k}}$ (it counts permutations and can be represented by either the $\sigma^{(z)}$-operator applied to one of the excited qubits, or by the $\sigma^{(x)}\sigma^{(z)}\sigma^{(x)}$-operator applied to one of the ground-state qubits (only to one!) of $A_{k}$).
-To arrange the exchange of states between $S_{j}$ and $S_{k}$, we introduce the projectors $P_{A_{k}}^{(j)}=|\tilde{j}\rangle_{A_{k}}\;{}_{A_{k}}\langle\tilde{j}|,$ $k=0,\dots N-2,$ $j=k+1,\dots N-1,$ $\tilde{j}=0,\dots N-k-2,$ and construct the operators $U_{S_{k}S_{j}}$:
+where $\sigma^{(x)}_{kj}$ includes the set of $\sigma^{(x)}$-operators which change the ground state of the ancilla $A_{k}$ to the state $|\tilde{j}\rangle_{A_{k}}$; $Z_{k}(j)$ changes the sign of the ancilla state $|\tilde{j}\rangle_{A_{k}}$ (it counts swaps and can be represented by either the $\sigma^{(z)}$-operator applied to one of the excited qubits, or by the $\sigma^{(x)}\sigma^{(z)}\sigma^{(x)}$-operator applied to one of the ground-state qubits (only to one!) of $A_{k}$).
+To arrange the exchange of states between $S_{j}$ and $S_{k}$, we introduce the projectors $P_{A_{k}}^{(j)}=|\tilde{j}\rangle_{A_{k}}\;{}_{A_{k}}\langle\tilde{j}|,$ $k=0,\dots N-2,$ $j=k+1,\dots N-1,$ $\tilde{j}=0,\dots N-k-2$, $\tilde{j}=j-k-1,$ and construct the operators $U_{S_{k}S_{j}}$:
 $U_{S_{k}S_{j}}$ $=$ $\Big{(}SWAP_{S_{k},S_{j}}\otimes P_{A_{k}}^{(j)}$ (12)
@@
 The circuits for the operators $V_{S_{j}A_{k}}$ and $U_{S_{k}S_{j}}^{(k)}$ are given in Fig.1(a) (the general matrix case) and in Fig.1(b) (the particular $4\times 4$-matrix case).
-Figure 1: $V$-and $U$-operators: (a) the general case; (b) the $4\times 4$-matrix case, then $\tilde{n}_{0}=2$, $\tilde{n}_{1}=\tilde{n}_{2}=1$, $\tilde{n}_{k}<n$, therefore, $A_{1}$ is a one-qubit ancilla.
-The depth of each operator $V_{S_{j}A_{k}}$ is $O(n)$ (due to the projectors $P_{S_{j}}^{(k)}$ with $n$ control qubits and we assume that all $\sigma^{(x)}$-operators of $\sigma^{(x)}_{kj}$ are applied in parallel to different qubits of $a_{k}$), while the depth of $U_{S_{k}S_{j}}$ is the of depth of the control-SWAP-operator which includes $n$ swaps each controlled by $n$ qubits, i.e.
-the total depth is $n\cdot n=n^{2}$ (or $O(n^{2})$).
-We prepare the operators
-$W_{k}=\prod_{j=k+1}^{N-1}U_{S_{k}S_{j}}V_{S_{j}A_{k}}.$ (13)
-Applying $W_{k}$ to the state $|\Phi_{0}\rangle|0\rangle_{A}$ we put $S_{k}$ into the state $|k\rangle_{S_{k}}$ in each term of the first part of (10).
+Figure 1: $V$ and $U$ operators for (a) the general case and (b) the $4\times 4$-matrix case.
+In the latter case, $\tilde{n}_{0}=\tilde{n}_{1}=2$, $\tilde{n}_{2}=1$, $\tilde{n}_{k}<n$.
+Hence, $A_{1}$ is a two-qubit ancilla.
+The depth of each operator $V_{S_{j}A_{k}}$ is $O(n)$ (due to the projectors $P_{S_{j}}^{(k)}$ with $n$ control qubits and we assume that all $\sigma^{(x)}$-operators of $\sigma^{(x)}_{kj}$ are applied in parallel to different qubits of $A_{k}$), while the depth of $U_{S_{k}S_{j}}$ is the depth of the control-SWAP-operator which includes $n$ swaps each controlled by $n$ qubits, i.e., the total depth is $nO(n)=O(n^{2})$.
+Theoretically, it can be implemented with less operations but this would not change the asymptotic complexity of the circuit.
+Applying the operators
+$W_{k}=\prod_{j=k+1}^{N-1}U_{S_{k}S_{j}}V_{S_{j}A_{k}},$ (13)
+to the state $|\Phi_{0}\rangle|0\rangle_{A}$, we put the subsystem $S_{k}$ into the state $|k\rangle_{S_{k}}$ in each term of the first part of (10).
 Although $W_{k}$ has product of $N-k-1$ pairs of $U$- and $V$-operators, only one pair of such operators is applied to each particular term in $|\Phi_{0}\rangle$ which is due to the structure of the projector operators $P^{(k)}_{S_{j}}$.
@@
 $|\Phi_{1}^{(D)}\rangle=W^{(1)}_{SA}|\Phi_{0}\rangle|0\rangle_{A}$ (16)
-$=\frac{\det(A)}{2^{\tilde{N}/2}}\;|0\rangle_{S_{0}}|1\rangle_{S_{1}}\otimes|{N-1}\rangle_{S_{N-1}}|0\rangle_{A}+|g_{2}\rangle_{SA},$
+$=\frac{\det(\tilde{A})}{2^{\tilde{N}/2}}\;|0\rangle_{S_{0}}|1\rangle_{S_{1}}\otimes|{N-1}\rangle_{S_{N-1}}|0\rangle_{A}+|g_{2}\rangle_{SA},$
 where
-$\det(A)$ $=$ $\sum_{k_{0}\neq\dots\neq k_{N-1}}\epsilon_{k_{0}\dots k_{N-1}}a_{k_{0}0}\dots a_{k_{N-1}(N-1)},$ (17)
+$\det(\tilde{A})$ $=$ $\sum_{k_{0}\neq\dots\neq k_{N-1}}\epsilon_{k_{0}\dots k_{N-1}}a_{k_{0}0}\dots a_{k_{N-1}(N-1)},$ (17)
 and $\epsilon_{k_{0}\dots k_{N-1}}$ is the permutation symbol.
 The depth of $W^{(1)}$ is $O(Nn^{2})$.
-Next, we can set states of all $S_{j}$ in the first term of $|\Phi_{1}\rangle$ into the $|0\rangle_{S_{j}}$ states.
+Next, we can set states of all $S_{j}$ in the first term of $|\Phi_{1}^{(D)}\rangle$ into the $|0\rangle_{S_{j}}$ states.
 To this end we apply the $\sigma^{x}$-operators to the excited qubits of $S$, i.e.
-prepare the operator $X_{S}$,
+prepare the operator $X_{S}$.
 $X_{S}=\prod_{j=1}^{N-1}X_{S_{j}}(j),$ (18)
-with $X_{S_{j}}(j)$ being the set of $\sigma^{(x)}$ operators acting on the excited qubits of $S_{j}$ in the state $|j\rangle_{S_{j}}$. The depth of $X_{S}$ is $O(1)$:
+with $X_{S_{j}}(j)$ being the set of $\sigma^{(x)}$ operators acting on the excited qubits of $S_{j}$ in the state $|j\rangle_{S_{j}}$.
+The depth of $X_{S}$ is $O(1)$.
+Applying $X_{S}$ to $|\Phi_{1}^{(D)}\rangle$ we obtain:
 $|\Phi_{2}^{(D)}\rangle=X_{S}|\Phi_{1}^{(D)}\rangle$ (19)
-$=\frac{\det(A)}{2^{\tilde{N}/2}}\;|0\rangle_{S}|0\rangle_{A}+|g_{3}\rangle_{SA}.$
-We notice that the operators $H_{A}$ in $W^{(1)}_{SA}$ and operator $X_{S}$ can be applied simultaneously because they effect on different qubits.
+$=\frac{\det(\tilde{A})}{2^{\tilde{N}/2}}\;|0\rangle_{S}|0\rangle_{A}+|g_{3}\rangle_{SA}.$
+We note that the operators $H_{A}$ in $W^{(1)}_{SA}$ and operator $X_{S}$ can be applied simultaneously because they affect different qubits.
 We separate them to underline the distinction in their functions.
@@
 $|\Phi_{3}^{(D)}\rangle=W^{(2)}_{SAB}|\Phi_{2}^{(D)}\rangle|0\rangle_{B}$ (21)
-$=\frac{\det(A)}{2^{\tilde{N}/2}}\;|0\rangle_{S}|0\rangle_{A}|1\rangle_{B}+|g_{3}\rangle_{SA}|0\rangle_{B}.$
-Now we measure the ancilla $B$ with the output $|1\rangle_{B}$. The probability of this result of measurement is $2^{-\tilde{N}}|{\det(A)}|^{2}$. After the measurement we get the following state:
-$|\Phi_{4}^{(D)}\rangle=|\Psi^{(D)}_{out}\rangle|0\rangle_{A},\;\;|\Psi^{(D)}_{out}\rangle=\arg(\det(A))\;|0\rangle_{S},$ (22)
+$=\frac{\det(\tilde{A})}{2^{\tilde{N}/2}}\;|0\rangle_{S}|0\rangle_{A}|1\rangle_{B}+|g_{3}\rangle_{SA}|0\rangle_{B}.$
+Now we measure the ancilla $B$ with the output $|1\rangle_{B}$.
+The probability of this result of measurement is $2^{-\tilde{N}}|{\det(\tilde{A})}|^{2}$.
+After the measurement we get the following state:
+$|\Phi_{4}^{(D)}\rangle=|\Psi^{(D)}_{out}\rangle|0\rangle_{A},\;\;|\Psi^{(D)}_{out}\rangle=\arg(\det(\tilde{A}))\;|0\rangle_{S},$ (22)
 The particular circuit for calculating the $4\times 4$ determinant using the proposed algorithm is given in Fig.
 2.
-The circuit for the determinant of $N\times N$ matrix includes $N\log N$ qubits of the subsystem $S$, $\tilde{N}=O(N\log N)$ qubits of the subsystem $A$ and one qubit of the ancilla $B$, therefore the size of this circuit is $O(N\log N)$.
-Figure 2: Quantum circuit illustrating computation of the determinant of a $4\times 4$ matrix.
-We omit superscript $(D)$ at $|\Phi_{k}\rangle$, $k=1,\dots,4$, and at $|\Psi_{out}\rangle$ for brevity.
-Operators $H_{A}$ and $X_{S}$ can be applied in parallel since they effect on different qubits.
+The circuit for the determinant of $N\times N$ matrix includes $N\log N$ qubits of the subsystem $S$, $\tilde{N}=O(N\log N)$ qubits of the subsystem $A$ and one qubit of the ancilla $B$, therefore the space of this circuit is $O(N\log N)$ qubits.
+Figure 2: Quantum circuit illustrating computation of determinant of a $4\times 4$ matrix.
+We omit superscript $(D)$ in $|\Phi_{k}\rangle$ $(k=1,\dots,4)$ and $|\Psi_{out}\rangle$ for brevity.
+Operators $H_{A}$ and $X_{S}$ can be applied in parallel since they affect different qubits.
 III Inverse matrix
-It is remarkable that the minor modification of the algorithm for calculating the determinant of a square $N\times N$ matrix in Sec.II allows to calculate the inverse of a square $(N-1)\times(N-1)$ matrix.
+We remark here that the minor modification of the algorithm for calculating the determinant of a square $N\times N$ matrix in Sec.II allows to calculate the inverse of a square $(N-1)\times(N-1)$ matrix.
 Moreover, this modification leads to a minor increase in the depth of the algorithm which becomes $O(N^{2}\log N)$ instead of $O(N\log^{2}N)$ in the algorithm for calculating the determinant.
-III.1 Preliminary arguments
+III.1 Preliminary consideration
 We impose a special structure on the elements in the first row and first column of the matrix $\tilde{A}$ given in (6):
-$a_{i0}=q,\;\;i=1,\dots,N-1,\;\;a_{0j}=1/\sqrt{N},\;\;j=0,\dots,N-1,$ (23)
+$a_{i0}=q,\;\;a_{0j}=1/\sqrt{N},\;\;i,j=0,\dots,N-1,$ (23)
 i.e.,
 $\tilde{A}=\left(\begin{array}[]{cccc}\frac{1}{\sqrt{N}}&\frac{1}{\sqrt{N}}&\cdots&\frac{1}{\sqrt{N}}\cr q&a_{11}&\cdots&a_{1(N-1)}\cr q&a_{21}&\cdots&a_{2(N-1)}\cr\cdots&\cdots&\cdots&\cdots\cr q&a_{(N-1)1}&\cdots&a_{(N-1)(N-1)}\end{array}\right).$ (29)
-The algorithm discussed below allows to invert the $(N-1)\times(N-1)$ matrix $A$,
-$A=\left(\begin{array}[]{ccc}a_{11}&\cdots&a_{1(N-1)}\cr a_{21}&\cdots&a_{2(N-1)}\cr\cdots&\cdots&\cdots\cr a_{(N-1)1}&\cdots&a_{(N-1)(N-1)}\end{array}\right),$ (34)
-(rather than the matrix $\tilde{A}$ itself) which is embedded into $\tilde{A}$.
+The algorithm discussed below allows to invert the $(N-1)\times(N-1)$ matrix $A$ (rather than the matrix $\tilde{A}$ itself) which is embedded into $\tilde{A}$,
+$A=\left(\begin{array}[]{ccc}a_{11}&\cdots&a_{1(N-1)}\cr a_{21}&\cdots&a_{2(N-1)}\cr\cdots&\cdots&\cdots\cr a_{(N-1)1}&\cdots&a_{(N-1)(N-1)}\end{array}\right).$ (34)
 According to the standard definition, the element $A^{-1}_{ji}$ of the inverse matrix is expressed in terms of the minor $M_{ij}$ obtained by removing the $i$th row and $j$th column of $A$:
@@
 $\frac{(-1)^{i+2}q}{\sqrt{N}}\left|\begin{array}[]{ccc}1&\cdots&1\cr a_{11}&\cdots&a_{1(N-1)}\cr\cdots&\cdots&\cdots\cr a_{(i-1)1}&\cdots&a_{(i-1)(N-1)}\cr a_{(i+1)1}&\cdots&a_{(i+1)(N-1)}\cr\cdots&\cdots&\cdots\cr a_{(N-1)1}&\cdots&a_{(N-1)(N-1)}\end{array}\right|,$ (43)
-where we take into account that $A_{ij}=a_{ij}$, $i,j=1,\dots,N-1$.
 In turn, this term includes the following term (remove the 1st row and $j$th column from the determinant (43); the $j$th column contains the elements $1$, $a_{1j}$, $a_{2j}$, …, $a_{(N-1)j}$):
-$-\frac{(-1)^{i+j}q}{\sqrt{N}}\left|\begin{array}[]{cccccc}a_{11}&\cdots&a_{1(j-1)}&a_{1(j+1)}&\cdots&a_{1(N-1)}\cr\cdots&\cdots&\cdots&\cdots&\cdots&\cdots\cr a_{(i-1)1}&\cdots&a_{(i-1)(j-1)}&a_{(i-1)(j+1)}&\cdots&a_{(i-1)(N-1)}\cr a_{(i+1)1}&\cdots&a_{(i+1)(j-1)}&a_{(i+1)(j+1)}&\cdots&a_{(i+1)(N-1)}\cr\cdots&\cdots&\cdots\cr a_{(N-1)1}&\cdots&a_{(N-1)(j-1)}&a_{(N-1)(j+1)}&\cdots&a_{(N-1)(N-1)}\end{array}\right|=$ (50)
-$-\frac{q}{\sqrt{N}}(-1)^{i+j}M_{ij}=-\frac{\det(A)q}{\sqrt{N}}A^{-1}_{ji}.$
+$-\frac{(-1)^{i+j}q}{\sqrt{N}}\left|\begin{array}[]{cccccc}a_{11}&\cdots&a_{1(j-1)}&a_{1(j+1)}&\cdots&a_{1(N-1)}\cr\cdots&\cdots&\cdots&\cdots&\cdots&\cdots\cr a_{(i-1)1}&\cdots&a_{(i-1)(j-1)}&a_{(i-1)(j+1)}&\cdots&a_{(i-1)(N-1)}\cr a_{(i+1)1}&\cdots&a_{(i+1)(j-1)}&a_{(i+1)(j+1)}&\cdots&a_{(i+1)(N-1)}\cr\cdots&\cdots&\cdots\cr a_{(N-1)1}&\cdots&a_{(N-1)(j-1)}&a_{(N-1)(j+1)}&\cdots&a_{(N-1)(N-1)}\end{array}\right|$ (50)
+$=-\frac{q}{\sqrt{N}}(-1)^{i+j}M_{ij}=-\frac{\det(A)q}{\sqrt{N}}A^{-1}_{ji}.$
 Thus, each element $A^{-1}_{ji}$ of the inverse matrix is included into $\det(\tilde{A})$ upto a factor which is the same for all elements.
-All what remains is to extract all the terms composing the element $A^{-1}_{ji}$ from those composing $\det(\tilde{A})$ and encode them into the proper state of some quantum subsystem.
-We show that this can be simply realized.
+The final step involves extracting all the terms that constitute the element $A^{-1}_{ji}$ from the terms that make up $\det(\tilde{A})$ and encode them into the appropriate state of some quantum subsystem.
+We show that this process can be easily realized.
 III.2 Basic algorithm
-The circuit for the algorithm is presented in Fig.3.
-Figure 3: Structure of a quantum circuit for matrix inversion. Ancillae are green on line.
+The circuit for the algorithm is presented in Fig.3(a).
 According to (7), the pure state $|\Phi_{0}\rangle$ of the whole system is given in (9) with normalization (8) and constraints (23) for the elements in the first row and first column of the matrix $\tilde{A}$, see (29).
-From this state, we select only those terms that compose the determinant of $\tilde{A}$ and therefore contain the terms that appear in the resulting inverse matrix $A^{-1}$:
-$|\Phi_{0}\rangle=\sum_{{k_{0},\dots,k_{N-1}=0}\atop{k_{0}\neq\dots\neq k_{N-1}}}^{N-1}a_{0k_{0}}\dots a_{(N-1)k_{N-1}}|k_{0}\rangle_{S_{0}}\cdots|k_{N-1}\rangle_{S_{N-1}}+|g_{1}\rangle_{S}$ (51)
-while the garbage $|g_{1}\rangle_{S}$ will be removed at the last step of the algorithm after ancilla measurement.
+From this state, we select only those terms that compose the determinant of $\tilde{A}$ and therefore contain the terms that appear in the resulting inverse matrix $A^{-1}$, i.e., we consider $|\Phi_{0}\rangle$ in form (10).
+Figure 3: (Color online) (a) Quantum circuit designed for matrix inversion (with the ancillae shown in green).
+(b) Quantum circuit designed for solving a system of linear algebraic equations.
+Both blocks of the circuit share the same ancillae (not shown in the accompanying figure).
 However, not all terms in the first part of $|\Phi_{0}\rangle$ are needed in our algorithm.
-We have to keep only those of them that include the 1st power of the factor $q$, (see constraints (23), i.e., exactly one state $|0\rangle_{S_{j}}$, $11\leq j\leq N-$ 1, can appear in the product $|k_{1}\rangle_{S_{1}}\dots|k_{N-1}\rangle_{S_{N-1}}$).
+We have to keep only those of them that include the 1st power of the factor $q$, (see constraints (23)), i.e., exactly one state $|0\rangle_{S_{j}}$, $11\leq j\leq N-$ 1, can appear in the product $|k_{1}\rangle_{S_{1}}\dots|k_{N-1}\rangle_{S_{N-1}}$.
 We also have to associate each of the probability amplitudes in (10) with the appropriate element $A^{-1}_{ij}$ of the inverse matrix.
@@
 The depth of $W_{SRC}^{(ij)}$ is $O(n)$.
-To establish this estimation we use the fact that the controlled operator $W_{SRC}^{(ij)}$ with $n$-qubit controlling register can be represented in terms of Toffoli gates by a circuit of depth $O(n)$ KShV and require that all $\sigma^{(x)}$ operators in $\sigma^{(x)}_{R^{(j)}C^{(i)}}$ can be applied simultaneously.
+To establish it we use the fact that the controlled operator $W_{SRC}^{(ij)}$ with $n$-qubit controlling register can be represented in terms of Toffoli gates by a circuit of depth $O(n)$ KShV and require that all $\sigma^{(x)}$ operators in $\sigma^{(x)}_{R^{(j)}C^{(i)}}$ can be applied simultaneously.
 Applying the operator ${\hat{W}}_{SRC}=\prod_{i=1}^{N-1}\prod_{j=1}^{N-1}W_{SRC}^{(ij)}$ (whose depth is $O(N^{2}n)$) to $|\Phi_{0}\rangle|0\rangle_{R}|0\rangle_{C}$ we obtain
-$|\Phi_{1}\rangle=\hat{W}_{SRC}|\Phi_{2}\rangle|0\rangle_{R}|0\rangle_{C}=$ (54)
-$\frac{1}{\sqrt{N}}\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}\sum_{{0\neq k_{1}\neq\dots\neq k_{i-1}}\atop{\neq k_{i+1}\neq\dots\neq k_{N-1}\neq j}}a_{1k_{1}}\dots a_{(i-1)k_{i-1}}a_{i0}a_{(i+1)k_{i+1}}\dots a_{(N-1)k_{N-1}}\times$
+$|\Phi_{1}\rangle=\hat{W}_{SRC}|\Phi_{2}\rangle|0\rangle_{R}|0\rangle_{C}$ (53)
+$=\frac{1}{\sqrt{N}}\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}\sum_{{0\neq k_{1}\neq\dots\neq k_{i-1}}\atop{\neq k_{i+1}\neq\dots\neq k_{N-1}\neq j}}a_{1k_{1}}\dots a_{(i-1)k_{i-1}}a_{i0}a_{(i+1)k_{i+1}}\dots a_{(N-1)k_{N-1}}\times$
 $|j\rangle_{S_{0}}|k_{1}\rangle_{S_{1}}\dots|k_{i-1}\rangle_{S_{i-1}}|0\rangle_{S_{i}}|k_{i+1}\rangle_{S_{i+1}}\dots|k_{N-1}\rangle_{S_{N-1}}|j\rangle_{R}|i\rangle_{C}+|g_{2}\rangle_{SRC},$
@@
 $|\Phi_{2}\rangle=W^{(1)}_{SA}|\Phi_{1}\rangle|0\rangle_{A}$ (54)
-$=-\frac{q\det(A)}{2^{(\tilde{N}+n)/2}}\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}A^{-1}_{ji}|j\rangle_{R}|i\rangle_{C}|0\rangle_{S_{0}}|1\rangle_{S_{1}}\otimes|{N-1}\rangle_{S_{N-1}}|0\rangle_{A}+|g_{1}^{(D)}\rangle_{SA}.$
+$=-\frac{q\det(A)}{2^{(\tilde{N}+n)/2}}\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}A^{-1}_{ji}|0\rangle_{S_{0}}|1\rangle_{S_{1}}$
+$\otimes|{N-1}\rangle_{S_{N-1}}|j\rangle_{R}|i\rangle_{C}|0\rangle_{A}+|g_{1}^{(D)}\rangle_{SA}.$
 Next, applying the operator $X_{S}=\prod_{j=1}^{N-1}X_{S_{j}}(j)$ given in (18) to $|\Phi_{2}\rangle$ we obtain
 $|\Phi_{3}\rangle=X_{S}|\Phi_{2}\rangle$ (55)
-$=-\frac{q\det(A)}{2^{(\tilde{N}+n)/2}}\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}A^{-1}_{ji}|j\rangle_{R}|i\rangle_{C}|0\rangle_{S}|0\rangle_{A}+|g_{2}^{(D)}\rangle_{SA},$
+$=-\frac{q\det(A)}{2^{(\tilde{N}+n)/2}}\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}A^{-1}_{ji}|0\rangle_{S}|j\rangle_{R}|i\rangle_{C}|0\rangle_{A}+|g_{2}^{(D)}\rangle_{SA},$
 Finally, applying $W^{(2)}_{SAB}$ (20) to $|\Phi_{3}\rangle|0\rangle_{B}$ we obtain
-$|\Phi_{4}\rangle=W^{(2)}_{SAB}|\Phi_{3}\rangle|0\rangle_{B}$ $=$ $\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}a_{ji}|j\rangle_{R}|i\rangle_{C}|0\rangle_{S}|0\rangle_{A}|1\rangle_{B}+|g\rangle_{SA}|0\rangle_{B},$ (57)
-$a_{ji}$ $=$ $-\frac{q\det(A)A^{-1}_{ji}}{2^{(\tilde{N}+n)/2}},$ (58)
+$|\Phi_{4}\rangle=W^{(2)}_{SAB}|\Phi_{3}\rangle|0\rangle_{B}$
+$=\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}a_{ji}|0\rangle_{S}|j\rangle_{R}|i\rangle_{C}|0\rangle_{A}|1\rangle_{B}+|g\rangle_{SA}|0\rangle_{B},$ (56)
+where
+$a_{ji}=-\frac{q\det(A)A^{-1}_{ji}}{2^{(\tilde{N}+n)/2}}.$ (57)
 The depth of the operator $D_{SAB}=W^{(2)}_{SAB}X_{S}W^{(1)}_{SA}$ (under certain assumptions on the physical realization of the operator $W_{k}$, see Eq.(13)) is $O(Nn^{2})<O(N^{2}n)$ (the depth of $\hat{W}_{SRC}$).
 After measuring the ancilla $B$ to get the output $|1\rangle$ with the probability $\frac{q^{2}|\det(A)|^{2}}{2^{\tilde{N}+n}}G^{2}\sim 2^{-O(\tilde{N})}=2^{-O(Nn)})$, $G=\sqrt{\sum_{i,j=1}^{N-1}|A^{-1}_{ij}|^{2}}$, we get the desired result
-$|\Phi_{5}\rangle=|\Psi_{out}\rangle|0\rangle_{S}|0\rangle_{A}.$ (59)
+$|\Phi_{5}\rangle=|\Psi_{out}\rangle|0\rangle_{S}|0\rangle_{A},$ (58)
 where
 $|\Psi_{out}\rangle$ $=$ $G^{-1}\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}A^{-1}_{ji}|j\rangle_{R}|i\rangle_{C}=G^{-1}|A^{-1}\rangle.$ (59)
-The state $|\Psi_{out}\rangle$ encodes $A^{-1}$ in the normalized form, i.e., $|\Psi_{out}\rangle\langle\Psi_{out}|=1$, while the state $|A^{-1}\rangle=\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}A^{-1}_{ji}|j\rangle_{R}|i\rangle_{C}$ is not normalized.
+The state $|\Psi_{out}\rangle$ encodes $A^{-1}$ in the normalized form, i.e., $\langle\Psi_{out}|\Psi_{out}\rangle=1$, while the state $|A^{-1}\rangle=\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}A^{-1}_{ji}|j\rangle_{R}|i\rangle_{C}$ is not normalized.
 The depth of discussed algorithm for matrix inversion is determined by the operator $\hat{W}_{SRC}$ and equals $O(N^{2}\log N)$.
-The circuit includes $N\log N$ qubits of the subsystem $S$, $\tilde{N}=O(N\log N)$ qubits of the ancilla $A$, one qubit of the ancilla $B$ and $2\log N$ qubits of $R$ and $C$, thus the size of this circuit approximately equals to the size of the circuit for calculating the determinant which is $O(N\log N)$.
-IV Application: solving systems of linear algebraic equations
-Having constructed the inverse matrix $A^{-1}$ encoded into the state of the subsystem $R\cup C$ we obtain a tool for solving a system of linear algebraic equations represented in the following matrix form:
-$A\bm{x}=\bm{b}.$ (61)
-Its solution reads
-$\bm{x}=A^{-1}\bm{b}.$ (62)
+The circuit includes $N\log N$ qubits of the subsystem $S$, $\tilde{N}=O(N\log N)$ qubits of the ancilla $A$, one qubit of the ancilla $B$ and $2\log N$ qubits of $R$ and $C$, thus the space of this circuit is approximated by the space of the circuit for calculating the determinant which is $O(N\log N)$.
+We note that this depth is obtained under assumption that the depth of the operator $W^{(1)}_{SA}$ in (15) is $N\log^{2}N$.
+If this depth is $N^{2}\log^{2}N$, then the depth of the matrix inversion algorithm is also $N^{2}\log^{2}N$.
+IV Solving systems of linear algebraic equations
+Having constructed the inverse matrix $A^{-1}$ encoded into the state of the subsystem $R\cup C$, we obtain a tool for solving a system of linear algebraic equations represented in the matrix form, $A\bm{x}=\bm{b}$, whose solution reads $\bm{x}=A^{-1}\bm{b}.$
 Therefore, we have to apply the multiplication algorithm proposed in Ref.ZQKW_arxiv2023 (see also Appendix, Sec.VI.1, for its modified version) to the matrices $A^{-1}$ and $\bm{b}$.
-The structure of the circuit is shown in Fig.4.
+The structure of the circuit is shown in Fig.3(b).
 More detailed circuit is presented in Fig.4.
-Figure 4: Structure of the circuit solving a system of linear algebraic equations. Both blocks share the same ancillae, which are not shown in figure.
-Notice that we can use the same ancillae in both blocks of this circuit (ancillae are not shown in Fig.4).
+Figure 4: (Color online) Circuit for quantum algorithm solving a system of linear algebraic equations $A\bm{x}=\bm{b}$.
+Circuit of matrix inversion (up to the operator $W^{(2)}_{SAB}$) is denoted by a block $A^{-1}$.
+All input qubits of the matrix-inversion algorithm (system $S$) become ancillae qubits.
+The output of the matrix-inversion algorithm (systems $R$ and $C$) together with system $b$ encoding the vector $\bm{b}$ become the input of the matrix-multiplication block.
+The output of the whole algorithm is encoded in the state $|x\rangle_{R}$ of the system $R$.
+All ancillae qubits are shown in green color.
+We omit the superscript $(L)$ in $|\Phi_{k}\rangle$ $(k=5,6,7)$ and $|\Psi_{out}\rangle$ for brevity.
+Notice that we can use the same ancillae in both blocks of this circuit (ancillae are not shown in Fig.3(b)).
 Moreover, the subsystem $S$ encoding the matrix $A$ can be used as ancilla in further calculations.
-In this case we do not measure ancillae $B$ (unlike Sec.III.2) and thus start with the state $|\Phi_{4}\rangle$ given in (57). The vector $\bm{b}$ is encoded as
+In this case we do not measure ancillae $B$ (unlike Sec.III.2) and thus start with the state $|\Phi_{4}\rangle$ given in (III.2). The vector $\bm{b}$ is encoded as
 $|b\rangle=\sum_{j=0}^{N-1}b_{j}|j\rangle_{b},\;\;b_{0}=0,\;\;\sum_{j=1}^{N-1}|b_{j}|^{2}=1.$ (60)
 The zero value for $b_{0}$ is prescribed by the encoding of $A^{-1}$ in (59), where the 00-states of $R$ and $C$ are not used.
-Out of system $S\cup A$ (which form ancilla in Fig.5), we take the one-qubit ancilla $\tilde{B}$ needed in the multiplication algorithm and neglect other qubits.
+Out of system $S\cup A$ (which forms an ancilla in Fig.4), we take the one-qubit ancilla $\tilde{B}$ needed in the multiplication algorithm and neglect other qubits.
 We also apply the operator $X_{B}=\sigma^{(x)}$ to invert the state of $B$.
 Thus, the state of the whole system reads
-$|\Phi^{(L)}_{5}\rangle$ $=$ $X_{B}|\Phi_{4}\rangle|b\rangle|0\rangle_{\tilde{B}}=a_{jl}b_{i}|j\rangle_{R}|l\rangle_{C}|i\rangle_{b}|0\rangle_{B}|0\rangle_{\tilde{B}}+|g_{1}\rangle_{RCb}|1\rangle_{B}|0\rangle_{\tilde{B}}.$ (64)
+$|\Phi^{(L)}_{5}\rangle=X_{B}|\Phi_{4}\rangle|b\rangle|0\rangle_{\tilde{B}}$
+$=a_{jl}b_{i}|j\rangle_{R}|l\rangle_{C}|i\rangle_{b}|0\rangle_{B}|0\rangle_{\tilde{B}}+|g_{1}\rangle_{RCb}|1\rangle_{B}|0\rangle_{\tilde{B}}.$ (61)
 Now we refer to the Appendix, Sec.VI.1, where the matrix multiplication algorithm is described.
@@
 the new operator reads:
-$W^{(3)}_{CBb\tilde{B}}=P_{CBb}\otimes\sigma^{(x)}_{\tilde{B}}+(I_{CBb}-P_{CBb})\otimes I_{\tilde{B}},\;\;P_{CBb}=|0\rangle_{C}|0\rangle_{B}|0\rangle_{b}\;{{}_{C}}\langle 0|_{B}\langle 0|{{}_{b}}\langle 0|.$ (65)
-Applying the operator $W^{(3)}_{CBb\tilde{B}}W^{(2)}_{C}W^{(1)}_{Cb}$ to $|\Phi^{(L)}_{5}\rangle$ (64) we obtain
+$P_{CBb}$ $=$ $|0\rangle_{C}|0\rangle_{B}|0\rangle_{b}\;{{}_{C}}\langle 0|_{B}\langle 0|{{}_{b}}\langle 0|,$
+$W^{(3)}_{CBb\tilde{B}}$ $=$ $P_{CBb}\otimes\sigma^{(x)}_{\tilde{B}}+(I_{CBb}-P_{CBb})\otimes I_{\tilde{B}}.$ (62)
+Applying the operator $W^{(3)}_{CBb\tilde{B}}W^{(2)}_{C}W^{(1)}_{Cb}$ to $|\Phi^{(L)}_{5}\rangle$ (IV) we obtain
 $|\Phi^{(L)}_{6}\rangle=W^{(3)}_{CBb\tilde{B}}W^{(2)}_{C}W^{(1)}_{Cb}|\Phi^{(L)}_{5}\rangle=$ (63)
@@
 The depth of the whole algorithm is defined by the depth of the subroutine inverting $A$ and thus equals $O(N^{2}n)$.
-In comparison with the circuit for inverse matrix, the circuit in Fig.
-5 includes $\log N$ qubits encoding the column $\bm{b}$ and thus remains essentially $O(N\log N)$.
-Figure 5: Circuit for the algorithm solving a system of linear algebraic equations (61).
-Circuit of matrix inversion (up to the operator $W^{(2)}_{SAB}$) is denoted by a block $A^{-1}$.
-All input qubits of the matrix-inversion algorithm (system $S$) become ancillae qubits.
-The output of the matrix-inversion algorithm (systems $R$ and $C$) together with system $b$ encoding the vector $\bm{b}$ become the input of the matrix-multiplication block.
-The output of the whole algorithm is encoded in the state $|x\rangle_{R}$ of the system $R$.
-All ancillae qubits are green on line.
-We omit the superscript $(L)$ at $|\Phi_{k}\rangle$, $k=5,6,7$ and at $|\Psi_{out}\rangle$ for brevity.
+In comparison with the circuit for inverse matrix, the circuit in Fig4 includes $\log N$ qubits encoding the column $\bm{b}$ and thus remains essentially $O(N\log N)$.
 V Conclusions
@@
 The size of the subsystem $S$ encoding the matrix $A$ is $N\log N$, the size of ancillae $A_{k}$, $k=0,\dots,N-2$, is $\tilde{N}=O(N\log N)$.
-Both these sizes serve to estimate the size of the whole algorithm which is $O(N\log N)$.
-It is remarkable that the $(N-1)\times(N-1)$-matrix inversion algorithm is essentially algorithm for calculating the determinant of the $N\times N$ matrix obtained by replacing the first row by the row of $1/\sqrt{N}$ and the first column by the column of $q$’s (up to the diagonal element).
+Both these sizes serve to estimate the space of the whole algorithm which is $O(N\log N)$ qubits.
+It is remarkable that the $(N-1)\times(N-1)$-matrix inversion algorithm is essentially the algorithm for calculating the determinant of the $N\times N$ matrix obtained by replacing the first row by the row of $1/\sqrt{N}$ and the first column by the column of $q$’s (up to the diagonal element).
 This special encoding provides the superposition of all products of matrix elements which are necessary for calculating minors included into the definition of the inverse matrix.
 The algorithm distributes these products with proper signs among the appropriate inverse-matrix elements which are encoded into the quantum state of certain subsystem thus forming the output state.
-The depth of the algorithm is $O(N^{2}\log N)$.
+The depth of the circuit is $O(N^{2}\log N)$.
 As an application of inversion we develop the algorithm for solving systems of linear algebraic equations.
 This algorithm consists of two blocks: matrix inversion and matrix multiplication.
-At that, the matrix inversion is most resource-consuming block which finally defines the depth of the algorithm (which is $O(N^{2}\log N)$), and probability of required ancilla state after measurement, $\sim 2^{-O(N\log N)}$.
-We have to notice that the depths of the algorithms for calculating the determinant and matrix inverse were obtained under assumption that the $U$-, $V$-operators in (13) can be applied in parallel.
+At that, the matrix inversion is most resource-consuming block which finally determines the depth of the algorithm (which is $O(N^{2}\log N)$), and probability of required ancilla state after measurement, $\sim 2^{-O(N\log N)}$.
+We have to note that the depths of the algorithms for calculating the determinant and matrix inverse were obtained under assumption that the $U$-, $V$-operators in (13) can be applied in parallel.
 Otherwise we have to change the depths $O(N\log^{2}N)$ and $O(N^{2}\log\,N)$ to $O(N^{2}\log^{2}N)$ in both cases.
 We emphasize that the matrix-multiplication algorithm is modified in comparison with that proposed in ZQKW_arxiv2023 .
-The detailed description of the modified version is presented in Appendix, Sec.
-VI.1.
+The detailed description of the modified version is presented in Appendix, Sec.VI.1.
 One notable aspect is the use of ancillae to gather all the garbage generated during the calculation, which can then be discarded with a measurement of a one-qubit ancilla.
-The weak point of both algorithms is probability of obtaining the required state of ancilla after measuring.
+The weak point of both algorithms is the probability of obtaining the required state of ancilla after measuring.
 This probability is $\sim 2^{-O(N\log N)}$ in all three algorithms.
 Combining the algorithm of matrix inversion with the algorithm of matrix multiplication we demonstrate the usage of two types of matrix encoding into the state of a quantum system.
-First one is the row-wise encoding used in the input of $N\times N$ matrix inversion algorithm (for simplicity, we assume that $N=2^{n}$).
+The first one is the row-wise encoding used in the input of algorithm for caculating the determinant of $N\times N$ matrix and matrix inversion algorithm (for simplicity, we assume that $N=2^{n}$).
 It requires $N$ subsystems of $\log N$ qubits each, and the state of $i$th subsystem encodes the elements of the $i$th row of the matrix.
@@
 The algorithm below is an alternative to that proposed in ZQKW_arxiv2023 .
-Both the depth and size of the algorithm are less than those parameters in ZQKW_arxiv2023 , although the estimations for them remains the same: depth is $O(\log K)$ and size is $O(\log K+\log N+\log M)$.
+Both the depth and size of the algorithm are less than those parameters in ZQKW_arxiv2023 , although the estimations for them remains the same: the depth is $O(\log K)$ and space is $O(\log K+\log N+\log M)$ qubits.
+Figure 5: (a) Circuit for the matrix multiplication algorithm. We omit superscript $(m)$ in $|\Phi_{k}\rangle$ $(k=0,\dots,4)$ and $|\Psi_{out}\rangle$ for brevity. (b) Notation for multiqubit CNOT.
 We first introduce one register of $n$ qubits, two registers of $k$ qubits and one register of $m$ qubits which enumerate rows and columns of both matrices.
@@
 However, if $C_{1}$ and $R_{2}$ are exchanged here and below in the description of the algorithm, it remains correct.
-All operators $W_{j}$, $j=1,\dots,k$ commute with each other.
-Applying the operator $W^{(1)}_{C_{1}R_{2}}$,
+The operator $W^{(1)}_{C_{1}R_{2}}$,
 $W^{(1)}_{C_{1}R_{2}}=\prod_{j=1}^{k}W_{j},$ (68)
-to $|\Phi_{0}^{(m)}\rangle$ we obtain:
+can be implemented by a depth-1 circuit, since all $W_{j}$ act on different qubits. Applying $W^{(1)}_{C_{1}R_{2}}$ to $|\Phi_{0}^{(m)}\rangle$ we obtain:
 $|\Phi_{1}^{(m)}\rangle=W^{(1)}_{C_{1}R_{2}}|\Phi_{0}^{(m)}\rangle=$ (69)
@@
 $|g_{1}^{(m)}\rangle_{R_{1}C_{1}R_{2}C_{2}},$
-selecting terms with $|0\rangle_{R_{2}}$ from the superposition state $|\Phi_{1}^{(m)}\rangle$ (the other terms are by definition garbage; their sum is denoted by $|g_{1}\rangle$ above).
+thus selecting terms with $|0\rangle_{R_{2}}$ from the superposition state $|\Phi_{1}^{(m)}\rangle$ (the other terms are by definition garbage; their sum is denoted by $|g_{1}\rangle$ above).
 Now we apply the Hadamard transformations
@@
 The circuit is given in Fig.
-6, its size is $12k+n+m+$ 1.
-Figure 6: (a) Circuit for the matrix multiplication algorithm; (b) Notation for multiqubit CNOT.
-We omit superscript $(m)$ at $|\Phi_{k}\rangle$, $k=0,\dots,4$, and at $|\Psi_{out}\rangle$ for brevity.
+5, its size is $12k+n+m+$ 1.